### v1

# spade: Synthesizing Assertions for Large Language Model Pipelines

Abstract.
Operationalizing large language models (LLMs) for custom, repetitive data pipelines is challenging, particularly due to their unpredictable and potentially catastrophic failures. Acknowledging the inevitability of these errors, we focus on identifying when LLMs may be generating incorrect responses when used repeatedly as part of data generation pipelines. We present spade, a method for automatically synthesizing assertions that identify bad LLM outputs. spade analyzes prompt version histories to create candidate assertion functions and then selects a minimal set that fulfills both coverage and accuracy requirements. In testing across nine different real-world LLM pipelines, spade efficiently reduces the number of assertions by 14% and decreases false failures by 21% when compared to simpler baselines.
1. Introduction
Developer Version History $\mathcal{P}_{\_}0,\ldots,\mathcal{P}_{\_}i$ All Examples Current Prompt Template $\mathcal{P}_{\_}{i+1}$ LLM Response $\spadesuit$ spadepteditsptformatsand sendsptsaves toptlabelsptgeneratesptsaves toptanalyzes $\Delta\mathcal{P}$s to generatecandidate assertionsptfilters assertionsbased on coverageand accuracy
Figure 1. Before a developer deploys a prompt template to production, spade analyzes the deltas (i.e., diffs) between consecutive prompt templates to generate assertions and uses a handful of labeled responses to filter out redundant and inaccurate assertions while maintaining coverage.
There is a lot of excitement around the use of large language models (LLMs) in data pipelines (Fernandez et al., 2023). Much of the enthusiasm is due to their simplicity: without needing large labeled datasets, one can easily create an automated, intelligent pipeline that executes any arbitrary data generation task within seconds—simply by prompting an LLM with natural language. However, operationalizing LLMs for unsupervised, repeated, or large-scale data generation tasks presents significant challenges (Parameswaran et al., 2023), such as potential errors, inappropriate responses, or hallucinations (Zamfirescu-Pereira et al., 2023; Shankar et al., 2022a).
Consider a movie streaming platform using LLMs to generate personalized movie recommendation notes. A developer might write a prompt template like: “write a personalized note for why a user should watch {movie_name} given the following information about the user: {personal_info} to be executed for multiple user-movie pairs. The templatized variables represent information that would get injected into the pipeline at runtime, allowing the pipeline to generate data for a variety of inputs. In theory, this prompt seems adequate, but the developer might observe some issues while testing it on a few example inputs: the LLM might reference a movie the user never watched, cite a sensitive attribute (e.g., race or ethnicity), or even have a basic issue like an overly short response.
To deal with such shortcomings, frameworks incorporating assertions into LLM pipelines have been proposed to filter out bad generations before reaching end-users (Rebedea et al., 2023; Singhvi et al., 2023). Given the unavoidable nature of LLM errors (Kalai and Vempala, 2023), such assertions are key to successful deployments of LLM pipelines that generate data. However, developers find it tough to write assertions for custom LLM pipelines (Parnin et al., 2023). Challenges include predicting all possible LLM failure modes, the time-consuming nature of writing assertions with various specification methods (like Python functions or LLM calls), the necessity for precision in assertions (especially those involving LLM calls), and the fact that many LLM pipeline developers lack software engineering expertise or coding experience (Kim et al., 2023; Zamfirescu-Pereira et al., 2023). Moreover, if there are too many or non-informative assertions, developers can be overwhelmed monitoring their results.
In this paper, we identify a new problem of automatically generating a good set of assertions for any data generation pipeline that leverages LLMs. While an assertion returns true (i.e., success) or false (i.e., failure) for an LLM response, a set of assertions returns the conjunction of individual assertions. A “good” set of assertions has minimal overlap and allows the developer to trade off between false successes and false failures. We decompose the problem into two components—candidate assertion generation and filtering—and present spade (System for Prompt Analysis and Delta-Based Evaluation, Figure 1). For candidate assertion generation, one may consider directly querying an LLM to “write assertions for x prompt,” but this may not cover the criteria the developer wants. We propose generating candidates from prompt deltas (i.e., diffs between two consecutive prompt versions), which often indicate specific failure modes of LLMs. For example, a developer might add an instruction like “avoid flowery language,” motivating an assertion to check the response language. We present an analysis of 19 custom data generation pipelines from users from LangChain (a startup that helps people build LLM pipelines) and pipeline prompt version histories, forming a taxonomy of prompt deltas. spade first automatically categorizes prompt deltas within the taxonomy, then synthesizes Python functions (that may include LLM calls) as candidate assertions. We demonstrate the potential of these assertions with a public release of this component of spade, with over 1300 uses across more than 10 sectors like finance, medicine, and IT (Shankar et al., 2023b).
Our analysis of candidate assertions revealed redundancy, inaccuracy, and a large number of functions, often exceeding 50 for just a few prompt deltas. Redundancy stems from repeated refinements to similar portions of a prompt or prompts with ambiguous instructions (e.g., “return a concise response”). Reducing this redundancy is not straightforward, even for engineers, since assertions can involve LLM calls with varying accuracies, motivating an automated filtering component. One approach is to use developer-labeled LLM responses to estimate each assertion’s false failure rate (FFR) and eliminate each one exceeding a developer-defined FFR threshold. However, the remaining assertions can have an FFR that cumulatively exceeds this threshold, and redundancies may persist. We show that selecting a small subset of assertions to meet failure coverage and FFR criteria is NP-hard. That said, we may express the problem as an integer linear program (ILP) and use an ILP solver to identify a solution. However, we find that labeled responses may not represent all failure modes, leading to omission of valuable assertions. For instance, in our movie recommendation scenario, an assertion that correctly verifies if the LLM-generated note is under 200 words will get discarded if all responses in our developer-labeled sample are below this limit. To expand coverage, active learning and weak supervision approaches can be used to sample and label new LLM prompt-response pairs for each candidate assertion (Castro and Nowak, 2008; Ratner et al., 2020), but this may be expensive or inaccessible for non-programmers. We introduce assertion subsumption as a way of ensuring comprehensive coverage: if one assertion doesn’t encompass the failure modes of another, both are selected. As such, spade selects a minimal set of assertions that respects failure coverage, accuracy, and subsumption constraints. Overall, we make the following contributions:
Prompt Deltas for Candidate Assertion Generation (Section 2). We identify prompt version history as a rich source of correctness criteria for data-generating LLM pipelines. We infer a taxonomy of assertion criteria based on prompt deltas from 19 LLM pipelines with version history, which we publish. We publicly released a tool [^1] https://spade-beta.streamlit.app to generate candidate assertions for any pipeline, providing early evidence for the utility of auto-generated assertions with over 1300 deployments while identifying opportunities to filter these assertions.
Filtering Method that Requires Little Data (Section 3). We show that selecting a minimal set of assertions while covering failure modes and meeting accuracy requirements is NP-hard. We express this problem as an ILP. Given that assertions may be redundant and inaccurate, and that LLM developers may not have engineering expertise or enough data to use existing data-driven approaches to ML pipeline testing (e.g., training models), we introduce assertion subsumption as a novel proxy for coverage in low-data settings.
Empirical Study (Section 4). We present spade, our system that automatically generates assertions for data-generating LLM pipelines. For nine LLM pipelines with prompt version history, we collect labeled prompt-response pairs (eight of which we open-source). spade  [^2] https://github.com/shreyashankar/spade-experiments generates assertions with good failure coverage and few false failures (i.e., failing a good response) across all pipelines. Our subsumption-based ILP outperforms simpler baselines that do not consider interactions between assertions by reducing assertions by 14% and lowering the false failure rate by 21%.
2. Identifying Candidate Assertions
Figure 2. Taxonomy of prompt deltas from 19 LLM pipelines.
Version $i$
$\Delta\mathcal{P}_{\_}i$
Possible New Assertion Criteria
1
+ Write a personalized note for why a user should watch {movie_name} given the following information about the user: {personal_info}.
Response should be personalized and relevant to the given user information
2
+ Include elements from the movie’s genre, cast, and themes that align with the user’s interests.
Response includes specific references to the user’s interests related to the movie’s genre, cast, and themes
3
+ Ensure the recommendation note is concise.
Response should be concise
4
- Ensure the recommendation note is concise. + Ensure the recommendation note is concise, not exceeding 100 words.
Response should be within the 100 word limit
5
- Include elements from the movie’s genre, cast, and themes that align with the user’s interests. + Mention the movie’s genre and any shared cast members between the {movie_name} and other movies the user has watched.
Response should mention genre and verify cast members are accurate
6
+ Mention any awards or critical acclaim received by movie_name.
Response should include references to awards or critical acclaim of the movie
7
+ Do not mention anything related to the user’s race, ethnicity, or any other sensitive attributes.
Response should not include references to sensitive personal attributes
Table 1. Comparison of 7 prompt versions for an LLM pipeline to write personalized movie recommendations. In each $\Delta\mathcal{P}_{\_}i$, a sentence starts with “+” if it is a newly added sentence; a sentence starts with “-” if it is removed from $\mathcal{P}_{\_}{i-1}$.
Our first goal is to generate a set of candidate assertions. We describe how prompt deltas can inform candidate assertions and explain how to derive candidate assertions from them.
2.1. Prompt Deltas
A single-step LLM pipeline consists of a prompt template $\mathcal{P}$, which is formatted with an input tuple $t$ to derive a prompt $p$, and returns a response $r$. There can be many versions of $\mathcal{P}$, depending on how a developer iterates on their prompt template. Let $\mathcal{P}_{\_}0$ be the empty string, the 0th version, and let $\mathcal{P}_{\_}i$ be the ith version of a template. In Section 1’s movie recommendation examples, suppose there are 7 versions, where $\mathcal{P}_{\_}7$ is the following: “Write a personalized note for why a user should watch {movie_name} given the following information about the user: {personal_info}. Ensure the recommendation note is concise, not exceeding 100 words. Mention the movie’s genre and any shared cast members between the {movie_name} and other movies the user has watched. Mention any awards or critical acclaim received by {movie_name}. Do not mention anything related to the user’s race, ethnicity, or any other sensitive attributes.”
We define a prompt delta $\Delta\mathcal{P}_{\_}{i+1}$ to be the diff between $\mathcal{P}_{\_}i$ and $\mathcal{P}_{\_}{i+1}$. Concretely, a prompt delta $\Delta\mathcal{P}$ is a set of sentences, where each sentence is tagged as an addition (i.e., “+”) or deletion (i.e., “-”). For example, Table 1 shows the $\Delta\mathcal{P}$s for a number of versions. Each sentence in $\Delta\mathcal{P}_{\_}i$ is composed of additions (i.e., new sentences in $\mathcal{P}_{\_}i$ that didn’t exist in $\mathcal{P}_{\_}{i-1}$) and deletions (i.e., sentences in $\mathcal{P}_{\_}{i-1}$ that don’t exist in $\mathcal{P}_{\_}{i}$). A modification to a sentence is represented by a deletion and addition—for example, $\Delta\mathcal{P}_{\_}{6}$ in Table 1 contains some new instructions added to a sentence from $\mathcal{P}_{\_}{5}$. Each addition in $\Delta\mathcal{P}_{\_}{i}$ indicates possible assertion criteria, as shown in the right-most column of Table 1.
2.2. Prompt Delta Analysis
We analyzed 19 LLM pipelines collected from LangChain users, each of which consists of between 3 and 11 historical prompt template versions. These pipelines span various tasks, from generating workout summaries to custom question-and-answering chatbots. Table 6 in Appendix B shows a summary of the pipelines, including a description of each pipeline and the number of prompt versions. For each pipeline, we categorized prompt deltas, i.e., $\Delta\mathcal{P}_{\_}i$, into different types—for example, instructing the LLM to include a new phrase in each response (i.e., inclusion), or instructing the LLM to respond with a certain tone (i.e., qualitative criteria). We iterated on our categories 4 times, ultimately producing the taxonomy in Figure 2. The taxonomy-annotated dataset of prompt versions can be found online [^3] https://github.com/shreyashankar/spade-experiments/blob/main/taxonomylabels.csv. Most prompt deltas fell under Data Integration (i.e., adding placeholder variables) and Workflow Description (i.e., adding instructions for the LLM to perform the task more accurately). In Table 2, we show sample prompt deltas for each category in our taxonomy, using the movie recommendation pipeline example from the introduction.
Category Explanation Example Prompt Delta
Response Format Instruction
Structure guidelines.
“+ Start response with ‘You might like…’.”
Example Demonstration
Illustrative example.
“+ For example, here is a response for sci-fi fans…”
Prompt Clarification
Refines prompt/removes ambiguity.
“- Discuss/+ Explain movie fit…”
Workflow Description
Describe “thinking” process.
“+ First, analyze viewing history…”
Data Integration
Adds placeholders.
“+ Include user’s {genre} reviews.”
Quantity Instruction
Adds numerical content.
“+ Keep note under 100 words.”
Inclusion Instruction
Directs specific content.
“+ Mention movie awards/acclaim.”
Exclusion Instruction
Advises on omissions.
“+ Avoid movie plot spoilers.”
Qualitative Criteria
Sets stylistic attributes.
“+ Maintain friendly, positive tone.”
Table 2. Categories of prompt deltas.
When deriving assertion candidates from automatically-identified prompt deltas, assertion quality clearly depends on the accuracy of the identified categories. Therefore, we confirmed GPT-4’s correct categorization of prompt deltas (as of October 2023). We assigned ground truth categories to all prompt versions from the 19 pipelines, and GPT-4 achieved an F1 score of 0.8135. The prompt used for category extraction from prompt deltas is detailed in Appendix C.
2.3. From Taxonomy to Assertions
Construct $\Delta\mathcal{P}_{\_}5$ “- Include elements from the movie’s genre, cast, and themes that align with the user’s interests. + Mention the movie’s genre and any shared cast members…” Prompt LLM with $\Delta\mathcal{P}_{\_}5$ to identify the delta type(s) and assertion criteria based on our taxonomy Synthesize Function(s)
Figure 3. Generating candidate assertion functions from a $\Delta\mathcal{P}$.
Using our taxonomy, spade employs an LLM to categorize and generate assertions for prompt deltas, detailed in Appendix C. For each $\Delta\mathcal{P}_{\_}i$, spade prompts the LLM to suggest as many criteria as possible for assertions, each aligning with a taxonomy category (e.g., Figure 3). A criterion is loosely defined as some natural language expression that operates on an example and evaluates to True or False (e.g., “check for conciseness”). spade analyzes every $\Delta\mathcal{P}_{\_}i$ instead of just last prompt version for several reasons: developers often remove instructions from prompts to reduce costs while expecting the same behavior (Parnin et al., 2023), prompts contain inherent ambiguities and imply multiple ways of evaluating some criteria, and complex prompts may lead to missed assertions if only one version is analyzed. Consequently, analyzing each $\Delta\mathcal{P}_{\_}i$ increases the likelihood of generating relevant assertions.
For each delta, spade collects the criteria identified and prompts the LLM again to create Python assertion functions. The synthesized functions can use external Python libraries or pose binary queries to an LLM for complex criteria. For function synthesis, the LLM is instructed that if the criterion is vaguely specified or open to interpretation, such as “check for conciseness,” it can generate multiple functions that each evaluate the criterion. In this conciseness example, the LLM could return multiple functions—a function that splits the response into sentences and ensures that there are no more than, say, 3 sentences, a function that splits the response into words and ensures that there are no more than, say, 25 words, or a function that sends the response to an LLM and asks whether the response is concise. The overall outcome of this step is a multiset of candidate functions $F=\{f_{\_}1,\dots,f_{\_}m\}$.
We adopt a two-step process in our approach, as it has been demonstrated that breaking tasks into steps can enhance LLM accuracy (Wei et al., 2022; Wu et al., 2022b; Grunde-McLaughlin et al., 2023). However, it’s worth noting that with daily advancements in LLMs, a one-step process might already be feasible. Additionally, while our taxonomy guides LLM-generated assertions now, future LLMs may implicitly learn these categories through reinforcement learning from human feedback (Christiano et al., 2017). Nevertheless, knowing the taxonomy-based categories associated with candidate assertions may help in filtering them.
2.4. LangChain Deployment
To assess the potential of our auto-generated assertions, in the first week of November 2023, we released an early prototype of spade’s candidate assertions via a Streamlit application [^4] https://spade-beta.streamlit.app/. At the time of public release, the Streamlit application had a different set of prompts to generate assertions; these prompts generated relatively few assertions compared to the number of assertions generated in this paper’s version of spade. However, the taxonomy has remained the same. In the Streamlit app, a developer can either paste their prompt template that they want to generate assertions for, or they can point to their LangChain Hub prompt template (which automatically contains prompt version history via commits). The app then visualizes the identified taxonomy categories in the user’s prompt and displays the candidate assertions, as shown in the screenshot in Figure 4.
We found significant interest in auto-generated LLM pipeline assertions: there were over 1300 runs of the app for custom prompt templates (i.e., not the sample default prompt template in app) within two weeks. These runs span a variety of fields, including medicine, education, cooking, and finance (Figure 5). Users for 10 of the 1300+ runs gave an optional “thumbs up” feedback on the assertions generated for them, and users for 43 of these runs clicked the “download assertions” button, which downloads the candidate assertions as a Python file. We note that users can copy the code in the assertions instead of downloading the assertions as a Python file, and we were unable to measure the copy events. No users clicked the “thumbs down” button—which is not to say that the generated assertions were perfect. In fact, anecdotally, we have seen the candidate assertion quality improve over the last several months as GPT-4 continually improves. On average, 3.3 assertions were generated per run, with minimum 1 and maximum 10 assertions generated for a run. Most of the runs corresponded to a single prompt version. We also found a deployment of spade-generated assertions from a LangChain user who built a “chat-with-your-pdf” tool [^5] https://twitter.com/thcalafatidis/status/1728144652119769394: in their words, “When I saw it I didnt beleive it could work that well, but it really did and made the evaluation process fun and ez.”
Figure 4. Screenshot of an early version of the spade Streamlit application.
[…]
Figure 5. Prompts submitted to the spade Streamlit application span a variety of fields.
In LLM pipelines with numerous prompt versions, we observed two main patterns. First, prompt engineering often leads to many similar assertions, and redundancy can be a headache at deployment when a developer has to keep track of so many assertions. For instance, a pipeline to summarize lecture transcripts [^6] https://smith.langchain.com/hub/kirby/simple-lecture-summary had 14 prompt versions, with many edits localized to the paragraph providing instructions on titles, speakers, and dates, creating multiple overlapping assertions. 10 assertions assessed the summary’s thesis; two are as follows:
Second, many assertions may be incorrect, causing runtime errors or false failures. Assessing the accuracy of these assertions is challenging, particularly for complex or LLM-invoking ones, where even experienced developers might not be able to gauge their effectiveness without viewing the results of the functions on many examples. Since there can be 50+ assertions generated for just a handful of prompt versions (as demonstrated in Section 4), manually filtering them by eyeballing failure rates for each subset of assertions is impractical. Therefore, we adopt an automated approach for filtering.
3. Filtering Candidate Assertions
Here, we address the issue of redundant and incorrect assertions identified in Section 2.4, particularly in pipelines with numerous prompt versions. Filtering this candidate set not only improves efficiency when deploying the assertions to run in production, but it also reduces cognitive overhead for the developer.
3.1. Definitions
Consider $e_{\_}i$ as an example end-to-end execution (i.e., run) of an LLM pipeline on some input. Let $E$ be the set of all such example runs (this set is not provided upfront, as we will deal with shortly). We define an assertion function $f:E\rightarrow\{0,1\}$, where 11 11 indicates success and 00 0 indicates failure. Let $F^{\prime}=\{f_{\_}1,f_{\_}2,\ldots,f_{\_}k\}$ be a set of $k$ assertions. An example $e_{\_}i$ is deemed successful by this set if and only if it satisfies all assertions in $F^{\prime}$. Specifically:
$\hat{y}_{\_}i=\begin{cases}1&\text{if }f(e_{\_}i)=1\quad\forall f\in F^{\prime},\\ 0&\text{otherwise.}\end{cases}$
We denote $y_{\_}i\in\{0,1\}$ to represent whether the LLM chain developer considers $e_{\_}i$ to be a success (1) or failure (0). Given all $m$ candidate assertions $F={f_{\_}1,f_{\_}2,\ldots,f_{\_}m}$, the objective is to select $F^{\prime}\subseteq F$ such that $\hat{y}_{\_}i=y_{\_}i$ for most examples in $E$, with $F^{\prime}$ being as small as possible. This goal involves maximizing failure coverage and minimizing the false failure rate and selected function count, expressed as follows:
Definition 3.1 ().
Coverage for a set $F^{\prime}$ is the proportion of actual failures that are correctly identified by $F^{\prime}$, defined as:
$\text{Coverage}\left(F^{\prime}\right)=\frac{\sum_{\_}{i}\mathbb{I}\left[\hat{y}_{\_}i=0\land y_{\_}i=0\right]}{\sum_{\_}{i}\mathbb{I}\left[y_{\_}i=0\right]}$
Definition 3.2 ().
False Failure Rate (FFR) for a set $F^{\prime}$ is the fraction of examples that $F^{\prime}$ incorrectly evaluates as failures ($\hat{y}_{\_}i=0$) when they are actually successful ($y_{\_}i=1$), defined as:
$\text{FFR}\left(F^{\prime}\right)=\frac{\sum_{\_}{i}\mathbb{I}\left[\hat{y}_{\_}i=0\land y_{\_}i=1\right]}{\sum_{\_}{i}\mathbb{I}\left[y_{\_}i=1\right]}$
In both Definition 3.1 and Definition 3.2, $\hat{y}_{\_}i$ represents set $F^{\prime}$’s prediction for the $i$-th example, $y_{\_}i$ is the actual outcome determined by the LLM chain developer, while $\mathbb{I}$ represents the indicator function. In practice,  Definition 3.1 and Definition 3.2 are impossible to compute since the universe of examples $E$ is unknown. So, for now, we assume access only to a subset $E^{\prime}\subset E$ of labeled LLM responses, where $E^{\prime}$ is a manually provided set of example runs and may not contain all the types of failures the LLM pipeline could hypothetically observe—an issue we will deal with in Section 3.3. Thus, we replace Definition 3.1 and Definition 3.2 with $\text{Coverage}_{\_}{E^{\prime}}\left(F^{\prime}\right)$ and $\text{FFR}_{\_}{E^{\prime}}\left(F^{\prime}\right)$, omitting the subscript $E^{\prime}$ for brevity. Table 7 in Appendix A summarizes the notation used throughout this section.
3.2. Coverage Problem Formulation
Our goal is therefore to select a minimal set of assertions $F^{\prime}\subseteq F$ based on a sample $E^{\prime}=\{e_{\_}1,\ldots,e_{\_}n\}\subset E$. Formally:
$\displaystyle\text{minimize}\quad|F^{\prime}|$
[…]
The above problem may be infeasible with certain values of $\alpha$, for multiple reasons. To see this, consider the case where no assertion catches a failure while $\alpha=1$. We ignore this (and similar cases) for now and defer their discussion to Section 4. To expand out the above, we introduce a matrix $M$ (size $n\times m$) to track each assertion’s result on each example $e_{\_}i$, where $M_{\_}{ij}=1$ if $f_{\_}j(e_{\_}i)=1$ and $M_{\_}{ij}=0$ otherwise. We also define binary variables $x_{\_}j$ and $w_{\_}{ij}$ to represent whether an assertion is chosen and if it marks an example as a failure: $w_{\_}{ij}=\left(1-M_{\_}{ij}\right)\cdot x_{\_}j$, which is based on whether $f_{\_}j$ denotes $e_{\_}i$ as a failure and $f_{\_}j$ is included in $F^{\prime}$. We additionally introduce the binary variable $u_{\_}i$ to represent whether a failed example is covered by any selected assertion:
$\displaystyle u_{\_}i\leq\sum_{\_}{j=1}^{m}w_{\_}{ij},\quad\forall i\in\left[1,n\right]:y_{\_}i=0.$
Then, the coverage constraint can be written as:
_$\displaystyle\frac{\sum_{\_}{i:y_{\_}i=0}u_{\_}i}{\sum_{\_}i\mathbb{I}\left[y_{\_}i=0\right]}\geq\alpha.$
Now, we formulate the FFR constraint. Observe that FFR can be decomposed into the following:
(1) _$\displaystyle\frac{\sum_{\_}{i=1}^{n}y_{\_}i\cdot\max_{\_}j\left(w_{\_}{ij}\right)}{\sum_{\_}{i=1}^{n}\left[y_{\_}i=1\right]}\leq\tau.$
The product $w_{\_}{ij}=(1-M_{\_}{ij})\cdot x_{\_}j$ in the numerator is 1 if $f_{\_}j$ is included in $F^{\prime}$ and incorrectly marks a good example as a failure. The $\max$ function ensures that as long as there is some selected function $f_{\_}j$ that marks $e_{\_}i$ as a failure, the product is 1. Multiplying $y_{\_}i$ by the $\max$ result ensures that the numerator is incremented only when $y_{\_}i=1$, or when the example $e_{\_}i$ is actually not a failure. This numerator is thus equivalent to the numerator in Definition 3.2, i.e., $y_{\_}i=1\land\hat{y}_{\_}i=0$. The entire fraction is less than the threshold $\tau$, which represents the maximum allowed false failure rate across the examples. To formulate FFR according to (1), we introduce a new binary variable $z_{\_}i$, which defines whether $F^{\prime}$ denotes $e_{\_}i$ as a failure while $e_{\_}i$ is actually a successful example (i.e., false failure):
$\displaystyle z_{\_}i\geq y_{\_}i\cdot w_{\_}{ij},\quad\forall i\in\left[1,n\right];\,\forall j\in\left[1,m\right].$
Then, the FFR constraint is:
_$\displaystyle\frac{\sum_{\_}{i=1}^{n}z_{\_}i}{\sum_{\_}{i=1}^{n}\mathbb{I}\left[y_{\_}i=1\right]}\leq\tau.$
We can then state the problem of minimizing the number of assertions while meeting $E^{\prime}$ coverage and FFR constraints as an Integer Linear Program (ILP):
$\displaystyle\text{minimize}\quad\sum_{\_}{j=1}^{m}x_{\_}j$
[…]
$\displaystyle\quad u_{\_}i\leq\sum_{\_}{j=1}^{m}w_{\_}{ij},\quad\forall i\in\left[1,n\right]\ \text{where}\ y_{\_}i=0;\quad\frac{\sum_{\_}{i:y_{\_}i=0}u_{\_}i}{\sum_{\_}i\mathbb{I}\left[y_{\_}i=0\right]}\geq\alpha;$
$\displaystyle\quad z_{\_}i\geq y_{\_}i\cdot w_{\_}{ij},\quad\forall i\in\left[1,n\right],\ \forall j\in\left[1,m\right];\quad\frac{\sum_{\_}{i=1}^{n}z_{\_}i}{\sum_{\_}{i=1}^{n}\mathbb{I}\left[y_{\_}i=1\right]}\leq\tau.$
We refer to a solution for this ILP as $\textsc{spade}_{\_}{\text{cov}}$. Trivially, the problem is NP-hard for $\tau=0$ and $\alpha=1$, via a simple reduction from set cover, and is in NP, since it can be stated in ILP form.
3.3. Subsumption Problem Formulation
So far, we’ve assumed that the developer is willing to provide a comprehensive set of labeled example runs $E^{\prime}$, In settings where the developer is unwilling to do so, and where $E^{\prime}$ does not include all failure types in $E$, $\textsc{spade}_{\_}{\text{cov}}$ may overlook useful assertions in $F$ that only catch failures in $E\setminus E^{\prime}$—as shown empirically in Section 4. We initially considered using active learning (Castro and Nowak, 2008) to sample more LLM responses for each assertion and weak supervision to label the responses (Ratner et al., 2020). However, this approach can be costly with state-of-the-art LLMs, and it demands significant manual effort to balance failing and successful examples for each assertion, ensuring meaningful FFRs and avoiding the exclusion of assertions due to underrepresented failure types. For this setting, we additionally introduce subsumption. Assuming that all candidate assertion functions cover as many failure modes as possible, our goal is to pick $F^{\prime}\subseteq F$ such that assertions in $F\setminus F^{\prime}$ are subsumed by $F^{\prime}$. Formally, a set of functions $S$ subsumes some function $f$ if the conjunction of functions in $S$ logically implies the conjunction of functions in $S$ and $f$. That is,
Definition 3.3 ().
A set of functions $S\implies f$ if and only if $\forall e\in E,\exists s\in S\text{ such that }s(e)\implies f(e)$. In other words, if $S\implies f$, then $f$ catches no new failures that $S$ does not already catch.
3.3.1. ILP with Subsumption Constraints
We reformulate the problem with subsumption constraints. Let $G$ be the set of functions in $F\setminus F^{\prime}$ not subsumed by $F^{\prime}$:
$\displaystyle\text{minimize}\quad|F^{\prime}|+|G|$
[…]
To represent $G$, we introduce binary variables and a matrix $K$ to denote subsumption relationships. $K_{\_}{ij}=1$ if and only if ${f_{\_}i}\implies f_{\_}j$. We discuss how we construct $K$ in more detail in Section 3.3.2. Recall that $x_{\_}j$ represents whether $f_{\_}j$ is selected in $F^{\prime}$. For each function $f_{\_}j$, a binary variable $r_{\_}j$ indicates if it is subsumed by $F^{\prime}$.
$\displaystyle r_{\_}j\leq\sum_{\_}{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{m}\left(x_{\_}{i}\cdot K_{\_}{ij}\right),\quad\forall j\in\left[1,m\right].$
Then, $s_{\_}j$ denotes if $f_{\_}j$ neither in $F^{\prime}$ nor subsumed by any function in $F^{\prime}$:
$\displaystyle s_{\_}j\leq 1-x_{\_}j,\quad s_{\_}j\leq 1-r_{\_}j,$ $\displaystyle\forall j\in\left[1,m\right];$
$\displaystyle s_{\_}j\geq(1-x_{\_}j)+(1-r_{\_}j)-1=1-x_{\_}j-r_{\_}j,$ $\displaystyle\forall j\in\left[1,m\right].$
Our objective is to minimize the sum of the number of functions in $F^{\prime}$ and non-subsumed functions $G$. The ILP formulation then becomes (with changes highlighted in blue):
$\displaystyle\text{minimize}\quad\sum_{\_}{j=1}^{m}x_{\_}j+{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}\pgfsys@color@rgb@stroke{0}{0}{0.6875}\pgfsys@color@rgb@fill{0}{0}{0.6875}\sum_{\_}{j=1}^{m}s_{\_}j}$
[…]
$\displaystyle\quad u_{\_}i\leq\sum_{\_}{j=1}^{m}w_{\_}{ij},\quad\forall i\in\left[1,n\right]\ \text{where}\ y_{\_}i=0;\quad\frac{\sum_{\_}{i:y_{\_}i=0}u_{\_}i}{\sum_{\_}i\mathbb{I}\left[y_{\_}i=0\right]}\geq\alpha;$
$\displaystyle\quad z_{\_}i\geq y_{\_}i\cdot w_{\_}{ij},\quad\forall i\in\left[1,n\right],\forall j\in\left[1,m\right];\quad\frac{\sum_{\_}{i=1}^{n}z_{\_}i}{\sum_{\_}{i=1}^{n}\mathbb{I}\left[y_{\_}i=1\right]}\leq\tau;$
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}\pgfsys@color@rgb@stroke{0}{0}{0.6875}\pgfsys@color@rgb@fill{0}{0}{0.6875}r_{\_}j\leq\sum_{\_}{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{m}(x_{\_}{i}\cdot K_{\_}{ij}),\quad\forall j\in\left[1,m\right];}$
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}\pgfsys@color@rgb@stroke{0}{0}{0.6875}\pgfsys@color@rgb@fill{0}{0}{0.6875}s_{\_}j\leq 1-x_{\_}j,\quad s_{\_}j\leq 1-r_{\_}j,\quad\forall j\in\left[1,m\right];}$
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}\pgfsys@color@rgb@stroke{0}{0}{0.6875}\pgfsys@color@rgb@fill{0}{0}{0.6875}s_{\_}j\geq 1-x_{\_}j-r_{\_}j,\quad\forall j\in\left[1,m\right].}$
We call a solution to this ILP $\textsc{spade}_{\_}{\text{sub}}$. We maintain the coverage constraint because the subsumption approach alone does not inherently account for the distribution or the significance of different types of failures. For instance, if a particular type of failure makes up a critical portion of $E^{\prime}$, a subsumption-based approach might overlook it. To see this in the simplest case, consider $\alpha=1$: simply optimizing for the sum $|F^{\prime}|+|G|$ does not guarantee all failures in $E^{\prime}$ are covered. In practice, $\textsc{spade}_{\_}{\text{sub}}$ is less sensitive to $\alpha$ than $\textsc{spade}_{\_}{\text{cov}}$, as we will discuss further in Section 4.
3.3.2. Assessing Subsumption
Here, we detail how to construct $K$, our matrix representing subsumption relationships between pairs of functions $f_{\_}i,f_{\_}j$. For pure Python functions, one could use static analysis to determine subsumption. However, it becomes complex when dealing with assertions that include LLM calls or a mix of pure Python and LLM-invoking assertions. For these, spade employs GPT-4 to identify potential subsumptions $\{a\}\implies b$ for pairs of functions $a,b$. For any pipeline, there are only two calls to the LLM to determine all subsumptions: first, all assertion functions are combined into a single prompt for GPT-4, instructing it to list as many subsumption relationships as it can identify, then prompting it again to transform its response into a parse-able list of pairs $a\implies b$. Details of this LLM subsumption prompt are in Appendix C.
To maximize precision of $\implies$ relationships identified, we employ some heuristics. First, $E^{\prime}$ can filter subsumptions: for $f_{\_}i$ and $f_{\_}j$, observe that:
[…]
In other words, any set containing $f_{\_}i$ definitely does not subsume $f_{\_}j$ if $f_{\_}j$ flags a failure that $\{f_{\_}i\}$ does not. Next, we use the FFR threshold to skip evaluating subsumption. Observe that, for any set of assertions $S$ and $f\not\in S$,
$\displaystyle\max\left(\text{FFR}\left(S\right),\text{FFR}\left(\{f\}\right)\right)$ $\displaystyle\leq\text{FFR}\left(S\cup\{f\}\right),$
(2) $\displaystyle\text{FFR}\left(S\cup\{f\}\right)$ $\displaystyle\leq\text{FFR}\left(S\right)+\text{FFR}\left(\{f\}\right).$
As such, we need not evaluate $\left\{f_{\_}i\right\}\implies f_{\_}j$ if either $\text{FFR}\left(\{f_{\_}i\}\right)\geq\tau$ or $\text{FFR}\left(\{f_{\_}j\}\right)\geq\tau$. Lastly, we use transitivity of implication to further prune checks: if ${x}\implies y$ and ${y}\implies z$, then ${x}\implies z$ is also true.
4. Evaluation
Pipeline # Good Ex. # Bad Ex. # Prompt Ver. Data Generation Task Full Prompt Link
codereviews 60 16 8
Writing reviews of GitHub repo pull requests
homanp/github-code-reviews
emails 43 55 3
Creating SaaS user onboarding emails
gitmaxd/onboard-email
fashion 48 34 16
Suggesting outfit ideas for specific events
N/A
financea 48 52 5
Summarizing financial earnings call transcripts
casazza/maptemplate
lecturesummariesb 27 22 14
Summarizing lectures or talks, focusing on main points and critical insights
kirby/simple-lecture-summary
negotiation 27 19 8
Writing tailored negotiation strategies based on provided contracts and target prices
antoniogonc/strategy-report
sportroutine 19 31 3
Transforming workout video transcripts into structured exercise routines
aaalexlit/sport-routine-to-program
statsbot 39 31 3
Writing interactive discussions for any topic in statistics
anthonynolan/statistics-teacher
threads 50 56 4
Crafting concise, engaging Twitter threads for specific audiences and topics
flflo/summarization
Table 3. Description of data-generating LLM pipelines in our experiments. The fashion examples (and ground-truth indicators of whether the example response is good or bad) are provided by a startup that uses LangChain. All other examples are synthetically generated except examples for the finance and lecturesummaries pipelines, which are taken from Kaggle.
Pipeline # Candidate Assertions Method FFR Coverage on $E^{\prime}$ Frac Func. Selected Frac Excl. Func. not Subsumed
codereviews 44 $\textsc{spade}_{\_}{\text{base}}$ 0.117 1 0.456 (20) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0 0.625 0.045 (2) 0.409 (18)
$\textsc{spade}_{\_}{\text{sub}}$ 0.117 0.875 0.341 (15) 0 (0)
emails 24 $\textsc{spade}_{\_}{\text{base}}$ 0 1 0.5 (12) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0 1 0.0417 (1) 0.458 (11)
$\textsc{spade}_{\_}{\text{sub}}$ 0 1 0.458 (11) 0 (0)
fashion 106 $\textsc{spade}_{\_}{\text{base}}$ 0.878 0.971 0.632 (67) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0.245 0.6 0.028 (3) 0.321 (34)
$\textsc{spade}_{\_}{\text{sub}}$ 0.224 0.62 0.377 (40) 0 (0)
finance 47 $\textsc{spade}_{\_}{\text{base}}$ 0.667 1 0.787 (37) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0.229 0.673 0.085 (4) 0.553 (26)
$\textsc{spade}_{\_}{\text{sub}}$ 0.208 0.981 0.553 (26) 0 (0)
lecturesummaries 70 $\textsc{spade}_{\_}{\text{base}}$ 0.528 1 0.457 (32) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0.194 0.643 0.014 (1) 0.414 (29)
$\textsc{spade}_{\_}{\text{sub}}$ 0.194 1 0.343 (24) 0 (0)
negotiation 50 $\textsc{spade}_{\_}{\text{base}}$ 0.444 1 0.4 (20) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0.222 0.632 0.04 (2) 0.32 (16)
$\textsc{spade}_{\_}{\text{sub}}$ 0.185 1 0.34 (17) 0 (0)
sportroutine 26 $\textsc{spade}_{\_}{\text{base}}$ 0.211 1 0.538 (14) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0.211 0.774 0.077 (2) 0.462 (12)
$\textsc{spade}_{\_}{\text{sub}}$ 0 0.871 0.308 (8) 0 (0)
statsbot 15 $\textsc{spade}_{\_}{\text{base}}$ 0 1 0.467 (7) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0 0.935 0.133 (2) 0.333 (5)
$\textsc{spade}_{\_}{\text{sub}}$ 0 1 0.467 (7) 0 (0)
threads 34 $\textsc{spade}_{\_}{\text{base}}$ 0 1 0.765 (26) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0 0.875 0.029 (1) 0.735 (25)
$\textsc{spade}_{\_}{\text{sub}}$ 0 1 0.589 (20) 0 (0)
Table 4. Results of different versions of spade with $\alpha=0.6$ and $\tau=0.25$. The check and x-marks denote whether $\alpha$ and $\tau$ constraints are met. Each entry is a fraction of the total number of candidate assertions for that pipeline (with the absolute number in parentheses). $\textsc{spade}_{\_}{\text{cov}}$ selects the fewest assertions overall. $\textsc{spade}_{\_}{\text{sub}}$ selects the fewest assertions while optimizing for subsumption.
We first discuss the LLM pipelines and datasets (i.e., $E^{\prime}$); then, we discuss methods and metrics and present our results. The experiment code, datasets, and LLM responses are hosted on GitHub [^7] https://github.com/shreyashankar/spade-experiments.
4.1. Pipeline and Dataset Descriptions
We evaluate spade on nine LLM pipelines—eight from LangChain Hub [^8] https://smith.langchain.com/hub, an open-source collection of LLM pipelines, and 1 proprietary pipeline. Six LangChain Hub pipelines helped develop the prompt delta taxonomy (Section 2.2), but two pipelines were added from spade’s Streamlit deployment (Section 2.4) after the taxonomy was created. The proprietary pipeline is the fashion pipeline, which suggests outfits for events. This pipeline is included due to its use of data not in LLM training and its real-world deployment—demonstrating spade’s practical applicability.
While we used real user prompt templates and histories (between 3 and 16 prompt versions), we constructed our own sets of examples prompt-response pairs and labels ($E^{\prime}$) for testing. Two datasets for the LangChain Hub pipelines were sourced from Kaggle, while the others were synthetically generated using Chat GPT Pro (based on GPT-4). For instance, for the codereviews pipeline that uses an LLM to review pull requests, we asked Chat GPT to generate placeholder values covering a variety of programming languages, application types, and diff sizes. We labeled the LLM responses for the 8 LangChain pipelines to assess whether they met the prompt instructions. The responses were split between GPT-3.5-Turbo and GPT-4. For the fashion pipeline, labeling was done by a developer at the corresponding startup. Table 3 provides details on each LLM pipeline and dataset. We have open-sourced all data for the 8 LangChain Hub pipelines.
4.2. Method Comparison and Metrics
As before, let $E^{\prime}$ be a dataset of example prompt-response pairs, as well as the corresponding labels of whether the response was good (i.e., 1) or bad (i.e., 0). Let $\tau$ be the FFR threshold and $F$ be the set of candidate assertions produced by the first step of spade (Section 2). If a candidate function $f$ results in a runtime error for some example $e$, we denote $f\left(e\right)=0$ (i.e., failure). All of our code was written in Python, using the PuLP Python package to find solutions for the ILPs. We used the default PuLP configuration, which uses the CBC solver (Forrest and Lougee-Heimer, 2005). We evaluated three versions of spade:
  • $\textsc{spade}_{\_}\text{base}$ selects all functions $f$ in $F$ where $\text{FFR}\left(\{f\}\right)\leq\tau$
  • $\textsc{spade}_{\_}\text{cov}$ is a solution to the ILP defined in Section 3.2
  • $\textsc{spade}_{\_}\text{sub}$ is a solution to the ILP defined in Section 3.3.1
Let $F^{\prime}$ represent the set of selected assertions by any version of spade. We measure four metrics:
  1. (1)
    Fraction of Assertions Selected (i.e., $|F^{\prime}|/|F|$)
  2. (2)
[…]
  3. (3)
    False Failure Rate (Definition 3.2)
  4. (4)
    Coverage on $E^{\prime}$ (Definition 3.1)
Additionally, an important aspect of $\textsc{spade}_{\_}{\text{sub}}$’s success is the effectiveness of subsumption assessment between all pairs of assertions. Since we do not have ground truth for subsumption, we focus on precision, calculated as the proportion of correctly identified subsumed pairs out of all subsumed pairs identified by the LLM. We do not assess recall—whether GPT-4 identified every possible subsumption—due to the impracticality of labeling possibly tens of thousands of assertion pairs per pipeline. Moreover, precision is more critical than recall or accuracy, as identifying even some subsumptions allows $\textsc{spade}_{\_}{\text{sub}}$ to achieve a solution with fewer selected assertions than $\textsc{spade}_{\_}{\text{base}}$.
4.3. Results and Discussion
Pipeline Precision
codereviews 0.90
emails 0.79
fashion 0.74
finance 0.79
lecturesummaries 0.89
negotiation 0.68
sportroutine 0.89
statsbot 0.86
threads 0.80
Table 5. Precision of assessing subsumption with GPT-4. Two authors verified the results.
Using GPT-4 to assess subsumption results in an average precision of 0.82 across all pipelines, as seen in Table 5, confirming its effectiveness. For simplicity, we set the coverage and FFR thresholds to be the same across all pipelines ($\alpha=0.6$, $\tau=0.25$). We report results for the three methods in Table 4. Consider the codereviews pipeline, for example, which uses an LLM to review a pull request for any code repository. Here, $\textsc{spade}_{\_}{\text{base}}$ selects 20 assertions, $\textsc{spade}_{\_}{\text{base}}$ selects two assertions, and $\textsc{spade}_{\_}{\text{sub}}$ selects 15 assertions. By selecting more functions, $\textsc{spade}_{\_}{\text{sub}}$ ensures that all non-subsumed functions are included. All three approaches respect the $E^{\prime}$ coverage constraint, but $\textsc{spade}_{\_}{\text{base}}$ violates the FFR constraint in 4 out of 9 pipelines.
On average, $\textsc{spade}_{\_}{\text{sub}}$ opts for approximately 14% fewer assertions compared to $\textsc{spade}_{\_}{\text{base}}$ and shows a significantly lower FFR, reducing it by about 21% relative to $\textsc{spade}_{\_}{\text{base}}$. $\textsc{spade}_{\_}{\text{cov}}$ excludes, on average, about 44% of functions that are not subsumed by $F^{\prime}$. We subsequently discuss the trade-offs between different spade implementations.
Subsumption vs. $E^{\prime}$ Coverage. $\textsc{spade}_{\_}{\text{cov}}$ and $\textsc{spade}_{\_}{\text{sub}}$ are complementary, the former being more useful if $E^{\prime}$ is more comprehensive. For our datasets, $E^{\prime}$ is definitely not comprehensive: Table 4 reveals that, on average, 44% of functions excluded by $\textsc{spade}_{\_}{\text{cov}}$ are not subsumed by the selected functions, despite being accurate within the FFR threshold. This is unsurprising given that each task has only 34 bad (i.e., failure) examples on average. While larger or more mature organizations may have extensive datasets and could get a meaningful result from $\textsc{spade}_{\_}{\text{cov}}$, $\textsc{spade}_{\_}{\text{sub}}$’s ability to select assertions that cover unrepresented potential failures can be beneficial in data-scarce settings. For example, here is a sample of 3 assertions for the codereviews pipeline ignored by $\textsc{spade}_{\_}{\text{cov}}$ but included in $\textsc{spade}_{\_}{\text{sub}}$ (with comments excluded for brevity):
Subsumption as a Means for Reducing Redundancy. Several pipelines exhibit a large discrepancy between functions selected in $\textsc{spade}_{\_}{\text{base}}$ and $\textsc{spade}_{\_}{\text{sub}}$, which occurs when there are many redundant candidates. For example, in the codereviews pipeline’s 8 prompt versions, the developer iterated several times on the instruction to give a clear and concise review, resulting in five assertions that check the same thing (two of which are shown below): Since all the five assertions meet the FFR constraint, individually, $\textsc{spade}_{\_}{\text{base}}$ would select them all, which is undesirable because they all do the same thing, but $\textsc{spade}_{\_}{\text{sub}}$ would select the one most compatible with the FFR constraint, as long as subsumption is assessed correctly. On the flip side, while assessing subsumption, the LLM may not recall all subsumptions, so $\textsc{spade}_{\_}{\text{sub}}$ may have duplicate assertions. For example, the codereviews pipeline contains assertions titled assert_includes_code_improvementv1 and assert_includes_code_improvementv2. $\boldsymbol{\alpha}$ and $\boldsymbol{\tau}$ Threshold Sensitivity. The feasibility of solutions from the ILP solver in spade is dependent on the chosen $\alpha$ and $\tau$ thresholds. If a feasible solution is not found, developers may need to adjust these values in a binary search fashion. In our case, all 9 LLM pipelines yielded feasible solutions with $\alpha=0.6$ and $\tau=0.25$. However, the small size of $E^{\prime}$ makes $\textsc{spade}_{\_}{\text{cov}}$ particularly sensitive to $\alpha$. In the pipelines, we observed that between one and five assertions covered 60% of $E^{\prime}$’s failures. For example, $\textsc{spade}_{\_}{\text{cov}}$ selected only one assertion for the emails pipeline: If $E^{\prime}$ is exhaustive of failure modes and representative of the distribution of failures (e.g., for the emails pipeline, most failures are actually due to the response lacking an encouragement to contact the company), $\textsc{spade}_{\_}{\text{cov}}$ might be a satisfactory solution. However, our $E^{\prime}$ datasets clearly were not exhaustive, considering that $\textsc{spade}_{\_}{\text{sub}}$ always chose additional assertions. $\textsc{spade}_{\_}{\text{sub}}$ is less sensitive to $\alpha$, as it explicitly selects assertions based on their potential to cover new failures (i.e., subsumption) without exceeding the FFR, even if the constraint on coverage is no longer tight. FFR Tradeoffs. Considering that the difference between the fraction of functions selected for $\textsc{spade}_{\_}{\text{base}}$ and $\textsc{spade}_{\_}{\text{sub}}$ is less than 10% for three LLM pipelines, one may wonder if the complexity of $\textsc{spade}_{\_}{\text{sub}}$ is worth it. $\textsc{spade}_{\_}{\text{sub}}$ is generally preferable because $\textsc{spade}_{\_}{\text{base}}$ fails to consistently meet the FFR threshold $\tau$. We observed that as prompt versions increase, so do the number of assertions, impacting $\textsc{spade}_{\_}{\text{base}}$ adversely. The worst-case FFR of a set is the sum of individual FFRs, as shown in Equation 2. Hence, with a large number of independent assertions, the total FFR is likely to surpass the threshold. This issue is evident in the fashion and lecturesummaries pipelines, where despite each of the 67 and 32 assertions meeting FFR constraints individually, the total FFR for $\textsc{spade}_{\_}{\text{base}}$ reaches 0.88 and 0.53, respectively. In practice, if spade were to be deployed in an interactive system, where spade could observe each LLM call in real-time (e.g., as a wrapper around the OpenAI API), the multitude of prompt versions further necessitates filtering assertions based on overall FFR. This underscores the need for the more complex $\textsc{spade}_{\_}{\text{cov}}$ or $\textsc{spade}_{\_}{\text{sub}}$ approaches.
4.4. Limitations and Future Work
Improving Quality of LLMs. While LLMs (both closed and open-source) are improving rapidly, and we don’t explicitly study prompt engineering strategies for spade, a complementary research idea is to explore such strategies or fine-tune small open-source models to generate assertions. Moreover, we proposed subsumption as a coverage proxy but didn’t explore prompt engineering or even non-LLM strategies (e.g., assertion provenance) for assessing subsumption. Despite LLM advancements, spade’s filtering stage remains crucial for reducing redundancy and ensuring accuracy, especially since assertions may involve LLMs.
Collecting Labeled Examples. Acquiring labeled data ($E^{\prime}$) is hard. Most of our datasets had very few prompt versions (only the ones committed to LangChain Hub), but in reality, developers may iterate on their prompt tens or hundreds of times. Future work could involve passive example collection via LLM API wrappers or gathering developer feedback on assertions. Prioritizing different types of assertions and formalizing these priorities within spade is another area for exploration. Additionally, assessing the accuracy of FFR estimates with limited $E^{\prime}$ and exploring methods to enhance FFR accuracy in the absence of large labeled datasets (e.g., via prediction-powered inference (Angelopoulos et al., 2023)), presents an interesting area for future work.
Supporting More Complex LLM Pipelines. Our study focuses on single-prompt LLM pipelines, but more complex pipelines, such as those involving multiple prompts, external resources, and human interaction, present opportunities for auto-generating assertions for each pipeline component. For instance, in retrieval-augmented generation pipelines(Lewis et al., 2020), assertions could be applied to the retrieved context before LLM responses are even generated.
5. Related Work
We survey work from prompt engineering, evaluating ML and LLMs, LLMs for software testing, and testing ML pipelines.
Prompt Engineering. For both nontechnical (Zamfirescu-Pereira et al., 2023) and technical users (Si et al., 2022; Parameswaran et al., 2023), prompt engineering is hard for several reasons: small changes in prompt phrasing (Liu et al., 2023; Arora et al., 2022) or the order of instructions or contexts (Lu et al., 2021) can significantly affect outputs. Moreover, as LLMs change under the hood of the API (i.e., prompt drift), outputs can change without developer awareness (Chen et al., 2023). Tools and papers are emerging to aid in prompt management and experimentation, and are even using LLMs to write prompts (Arawjo et al., 2023; Wu et al., 2022b, a; Cheng et al., 2023; Zhou et al., 2022). Moreover, deployed prompts introduce new challenges, like “balancing more context with fewer tokens” and “wrangling prompt output” to meet user-defined criteria (Parnin et al., 2023). Our work doesn’t focus explicitly on helping developers create better prompts, but it could indirectly support developers in improving prompts through recommended assertions.
ML and LLM Evaluation. Evaluating and monitoring deployed ML models is known to be challenging (Paleyes et al., 2022; Shankar et al., 2022b). Evaluating LLMs in a deployed setting is even more challenging because LLMs are typically used for generative tasks, where the outputs are free-form (Dou et al., 2021). Some LLM pipeline types, like question-answering with retrieval-augmented generation pipelines (Lewis et al., 2020), can use standardized, automated metrics (Es et al., 2023; Saad-Falcon et al., 2023), but others face challenges due to unknown metrics and lack of labeled datasets (Xiao et al., 2023; Chang et al., 2023; Parnin et al., 2023). Typically, organizations rely on human evaluators for LLM outputs (Gehrmann et al., 2023; Wang et al., 2023b; Parnin et al., 2023), but recent studies suggest LLMs can self-evaluate effectively with detailed “scorecards” (Chan et al., 2023; Zhang et al., 2023; Kim et al., 2023). However, writing these scorecards can be challenging (Parnin et al., 2023), motivating auto-generated evaluators.
LLMs for Software Testing. LLMs are increasingly being used in software testing, mainly for generating unit tests and test cases (Wang et al., 2023a; Schäfer et al., 2023; Siddiq et al., 2023; Steenhoek et al., 2023; Lemieux et al., 2023). Research explores how LLMs’ prompting strategies, hallucinations, and nondeterminism affect code or test accuracy (Ouyang et al., 2023; Feldt et al., 2023; Chen et al., 2021; Fan et al., 2023). Our work is complementary and leverages LLMs to generate code-based assertions for LLM pipelines.
Testing in ML Pipelines. ML pipelines are hard to manage in production. Much of the literature on ML testing is geared towards validating structured data, through analyzing data quality (Hynes et al., 2017; Breck et al., 2019; Schelter et al., 2018; Shankar et al., 2023a) or provenance (Schelter et al., 2023; Namaki et al., 2020). Platforms for ML testing typically offer automated experiment tracking and prevention against overfitting (Renggli et al., 2019; Agrawal et al., 2019), as well as tools for data distribution debugging (Grafberger et al., 2022). Model-specific assertions typically require human specification (Kang et al., 2020), or at least large amounts of data to train learned assertions (Kang et al., 2022). LLM chains or pipelines are a new class of ML pipelines, and LLMs themselves can generate assertions with little data. A recent study highlights the difficulty of testing LLM pipelines for “copilot”-like products: developers want to ensure accuracy while avoiding excessive resource use, such as running hundreds of assertions (Parnin et al., 2023)—motivating our assertion filtering approach.
6. Conclusion
Our work introduced a new problem of auto-generating assertions to catch failures in LLM pipelines, as well as a solution that comprises of two components: first, it synthesizes candidate assertions; then, it filters them. We proposed a taxonomy of prompt edits for assertion synthesis, demonstrating its potential via integration and deployment with LangChain. We expressed the selection of an optimal set of assertions for covering failures with high accuracy as an Integer Linear Program (ILP). We proposed assertion subsumption to cover failures in data-scarce scenarios and incorporated this into our ILP. Our auto-generating assertion system, spade, was evaluated on nine real-world data-generating LLM pipelines. We have made our code and datasets public for further research and analysis.
Appendix A Notation
Table 7 summarizes the important notation used in the paper, including high-level descriptions of the ILP variables.
Symbol Description
$\mathcal{P}$
A prompt template (i.e., a string with placeholders, intended to be submitted to an LLM)
$\mathcal{P}_{\_}i$
The $i$th version of a prompt template
$\Delta\mathcal{P}_{\_}i$
The diff or changes between consecutive versions of a prompt template
$E$
The set of all hypothetical example runs for an LLM pipeline
$E^{\prime}$
A small set of example runs for an LLM pipeline (with labels denoting whether responses were good) that we can observe, where $E^{\prime}\subset E$
$n$
The number of examples in $E^{\prime}$
$e_{\_}i$
The $i$th example in $E^{\prime}$
$f$
An arbitrary assertion function
$F$
The set of all candidate assertions from Section 2
$m$
The number of assertions in $F$
$F^{\prime}$
A subset of $F$ selected as a minimal set of assertions
$G$
Set of functions in $F\setminus F^{\prime}$ not subsumed by $F^{\prime}$ (as defined by Definition 3.3)
$\hat{y}_{\_}i$
Binary indicator of whether example $e_{\_}i$ satisfies all assertions in $F^{\prime}$
$y_{\_}i$
Binary indicator of whether a developer considers the LLM pipeline response for example $e_{\_}i$ good (i.e., success) or bad (i.e., failure)
$\alpha$
Threshold for coverage of failing examples; $F^{\prime}$ should cover at least $\alpha$ of failing examples
$\tau$
Threshold for False Failure Rate (FFR); $F^{\prime}$ should fail no more than $\tau$ good examples
$M$
$n\times m$ matrix representing the results of assertions on examples in $E^{\prime}$
$K$
$m\times m$ matrix representing subsumption relationships between functions
$x_{\_}j$
ILP variable indicating inclusion of $f_{\_}j$ in $F^{\prime}$
$w_{\_}{ij}$
ILP variable representing if $F^{\prime}$ denotes $e_{\_}i$ as a failure
$u_{\_}i$
ILP variable indicating if a failed example is covered
$z_{\_}i$
ILP variable defining if $F^{\prime}$ incorrectly marks a successful example as a failure (false failure)
$r_{\_}j$
ILP variable indicating if a function is subsumed by any function in $F^{\prime}$
$s_{\_}j$
ILP variable representing functions in $F\setminus F^{\prime}$ not subsumed by $F^{\prime}$
Table 7. Notation used in the paper
Appendix B LangChain Prompt Hub and Taxonomy of Prompt Deltas
The LangChain Prompt Hub is an open-source repository of prompts for chains, detailing the version history of the prompts. In analyzing prompt deltas, we filter the Prompt Hub for user-uploaded chains with at least 3 prompt versions. Table 6 summarizes the chains analyzed.
Appendix C spade Prompts
spade leverages LLMs in three places. The first uses GPT-4 to categorize the delta (constructed by Python’s difflib). The second uses GPT-4 to generate Python assertion functions, based on the prompt and categories identified by the first ste.. The third usage of LLMs is in asking GPT-4 to evaluate whether two functions subsume each other.
Given prompt_diff, a list of sentences that have been modified from the previous prompt template to the current prompt template (i.e., $\Delta\mathcal{P}$), the prompt for categorizing the delta is as follows:
Let concepts be the parsed categories identified by the previous prompt. The prompt for generating the Python assertion functions is as follows:
C.1. Evaluating Subsumption
To evaluate subsumption, we use one prompt to query the subsumed pairs, given all assertions, and then a second prompt to format the subsumed pairs as a JSON so it can be easily parsed by spade. The first prompt is as follows:
In the prompt above, the assertion_blob represents a string of all assertion functions. Then, the second prompt is as follows: