### v2

# spade: Synthesizing Data Quality Assertions for Large Language Model Pipelines

Abstract.
Large language models (LLMs) are being increasingly deployed as part of pipelines that repeatedly process or generate data of some sort. However, a common barrier to deployment are the frequent and often unpredictable errors that plague LLMs. Acknowledging the inevitability of these errors, we propose data quality assertions to identify when LLMs may be making mistakes. We present spade, a method for automatically synthesizing data quality assertions that identify bad LLM outputs. We make the observation that developers often identify data quality issues during prototyping prior to deployment, and attempt to address them by adding instructions to the LLM prompt over time. spade therefore analyzes histories of prompt versions over time to create candidate assertion functions and then selects a minimal set that fulfills both coverage and accuracy requirements. In testing across nine different real-world LLM pipelines, spade efficiently reduces the number of assertions by 14% and decreases false failures by 21% when compared to simpler baselines. spade has been deployed as an offering within LangSmith, LangChain’s LLM pipeline hub, and has been used to generate data quality assertions for over 2000 pipelines across a spectrum of industries.
1. Introduction
$\spadesuit$ spade Developer Current Prompt Template $\mathcal{P}_{\_}{i+1}$ Version History $\mathcal{P}_{\_}0,\ldots,\mathcal{P}_{\_}i$ LLM Output All Example Inputs and Outputs 1 Generate Assertion Candidates 2 Filter Bad Assertions Assertions to Deploy editssaves tolabelssaves to
Figure 1. Before a developer deploys a prompt template to production, spade analyzes the deltas (i.e., diffs) between consecutive prompt templates to generate assertions. Then, spade uses labeled pipeline inputs and outputs to filter out redundant and inaccurate assertions, while maintaining coverage of bad outputs.
There is a lot of excitement around the use of large language models (LLMs) for processing, understanding, and generating data (Fernandez et al., 2023). Without needing large labeled datasets, one can easily create an LLM pipeline for any task on a collection of data items—simply by crafting a natural language prompt that instructs the LLM on what to do with each item. This could span traditional data processing tasks, such as summarizing each document in a corpus, extracting entities from each news article in a collection of articles, or even imputing missing data for each tuple in a relation (Narayan et al., 2022). LLMs additionally enable new and more complex data processing tasks that involve generating data, e.g., an LLM could write an explanation for why a product was recommended to a user, author emails to potential sales leads, or draft blog posts for marketing and outreach. In all of these data processing tasks, deploying these LLM pipelines at scale—either offline, on each batch of data items, or online, as and when new items arrive—presents significant challenges, due to data quality errors made by LLMs seemingly at random (Kalai and Vempala, 2023)—with LLMs often disregarding instructions, making mistakes with the output format, or hallucinating facts (Zamfirescu-Pereira et al., 2023; Shankar et al., 2022a).
One approach to catch errors in deployed LLM pipelines is via data quality assertions. Indeed, multiple recent papers (Rebedea et al., 2023; Singhvi et al., 2023; Kim et al., 2023) and LLM pipeline authoring systems (Guardrails, 2023; Langchain, 2023; Llama Index, 2023) provide mechanisms to embed manually-provided or selected assertions as part of LLM pipelines to catch errors during deployment. However, determining which assertions to add remains an open problem—and is a big customer painpoint based on our experience at LangChain—a company that helps people build LLM pipelines. Developers often find it difficult to determine the right set of assertions for their custom tasks (Parnin et al., 2023). Challenges include predicting all possible LLM failure modes, the time-consuming nature of writing assertions with various specification methods (like Python functions or LLM calls), the necessity for precision in assertions (especially those involving LLM calls), and the fact that many LLM pipeline developers lack software engineering expertise or coding experience (Kim et al., 2023; Zamfirescu-Pereira et al., 2023). Moreover, if there are non-informative assertions or too many of them, developers can get overwhelmed monitoring the results of these assertions. While there is some work on automatically detecting errors in traditional ML pipelines (Schelter et al., 2018; Shankar et al., 2023a; Polyzotis et al., 2018; Breck et al., 2019a), this line of work operates on many structured records at a time, and doesn’t apply to an unstructured setting. So, we target the following question: can we identify data quality assertions for LLM pipelines with as little effort from developers as possible?
Example LLM Pipeline. Consider an LLM pipeline for a movie streaming platform, where the task is to generate a paragraph of text explaining why a specific movie was recommended to a specific user. A developer might write a prompt template like: “Given the following information about the user, {personal_info}, and information about a movie, {movie_info}: write a personalized note for why the user should watch this movie” to be executed for many user-movie pairs. In theory, this prompt seems adequate, but the developer might observe some data quality issues while testing it across different inputs: the LLM output might reference a movie the user never watched, cite a sensitive attribute (e.g., race or ethnicity), or even exhibit a basic issue by being too short. To fix these problems, developers typically add instructions to the prompt to catch these data quality issues, such as “Don’t reference race in your answer”. However, the LLM may still violate these instructions in an unpredictable manner for some data items, requiring assertions applied to LLM outputs post-hoc, during deployment.
Analyzing Prompt Version Histories. To automatically synthesize assertions for developers, our first insight is that we can mine prompt version histories to identify assertion criteria for LLM pipelines, since developers implicitly embed data quality requirements through changes to the prompt, or prompt deltas, over time. In our example above, instructions such as “Make sure your response is at least 3 sentences” or “Don’t reference race in your answer” could each correspond to a candidate assertion. To understand the types of prompt deltas in LLM pipelines and verify their usefulness for data quality assertions, we present an analysis of prompt version histories of 19 custom pipelines from LangChain users. Using this analysis, we construct a taxonomy of prompt deltas (Figure 2), which may be of independent interest for researchers and practitioners studying how to best build LLM pipelines.
Redundancy in Data Quality Assertions. Our second insight from analyzing these pipelines is that if we were to create candidate assertions from prompt deltas, say, using an LLM, there may be far too many assertions—often exceeding 50 for just a few prompt deltas. Many of these assertions (or equivalently, prompt deltas) are redundant, while some are too imprecise or ambiguous to be useful (e.g., “return a concise response”). Reducing this redundancy is nontrivial, even for engineers: assertions themselves can involve LLM calls with varying accuracies, motivating an automated component that can filter out bad candidates. One approach is to use a small handful of developer-labeled LLM outputs to estimate each data quality assertion’s false positive and false negative rate—but collectively determining the right set of assertions that can catch most errors, while incorrectly flagging as few outputs as possible, is not straightforward.
spade: Our Data Quality Assertion Generation Framework. In this paper, we leverage the aforementioned insights to develop spade (System for Prompt Analysis and Delta-Based Evaluation, Figure 1). spade’s goal is to select a set of boolean assertions with minimal overlap, while maximizing coverage of bad outputs and minimizing false failures (correct outputs that are incorrectly flagged) for the conjunction of selected assertions. We decompose spade into two components—candidate assertion generation and filtering.
Component 1: Prompt Deltas for Candidate Assertion Generation. For generating candidate assertions, instead of directly querying an LLM to “write assertions for x prompt,” which causes the LLM to miss certain portions of the prompt, we generate candidates from each prompt delta, which typically indicate specific failure modes of LLMs. spade leverages the aforementioned taxonomy of prompt deltas we constructed by first automatically categorizing deltas using the taxonomy, then synthesizing Python functions (that may include LLM calls) as candidate assertions. At LangChain, we publicly release this component of spade—which has been subsequently used for over 2000 pipelines across more than 10 sectors like finance, medicine, and IT (Shankar et al., 2023b). We present an analysis of this usage in Section 2.4.
Component 2: Filtering Candidate Assertions with Limited Data. To filter out incorrect and redundant candidate assertions, instead of requiring cumbersome manual selection or even fine-tuning of separate models (Saad-Falcon et al., 2023; Wang et al., 2023b), we propose an automated approach that only requires a small handful of labeled examples, which are usually already present in most target applications. Using these examples, we could estimate each assertion’s false failure rate (FFR), i.e., how often an assertion incorrectly flags failures, and eliminate individual assertions that exceed a given threshold. However, given we are selecting a set of assertions, the set may still exceed the FFR threshold and flag too many failures incorrectly, and redundancies may persist. We show that selecting a small subset of assertions to meet failure coverage and FFR criteria is NP-hard. That said, we may express the problem as an integer linear program (ILP) and use an ILP solver to identify a solution in a reasonable time given the size of our problem (hundreds to thousands of variables).
In some cases, when there are limited developer-provided examples, we find that labeled LLM outputs may not cover all failure modes, leading to omission of valuable data quality assertions. For instance, in our movie recommendation scenario, an assertion that correctly verifies if the output is under 200 words will get discarded if all outputs in our developer-labeled sample respect this limit. To expand coverage, active learning and weak supervision approaches can be used to sample and label new LLM input-output pairs for each candidate assertion (Castro and Nowak, 2008; Ratner et al., 2020), but this may be expensive or inaccessible for non-programmers. We introduce assertion subsumption as a way of ensuring comprehensive coverage: if one assertion doesn’t encompass the failure modes of another, both may be selected. As such, spade selects a minimal set of assertions that respects failure coverage, accuracy, and subsumption constraints.
Overall, we make the following contributions:
  • We identify prompt version history as a rich source for LLM output correctness, creating a taxonomy of assertion criteria from 19 diverse LLM pipelines (Section 2),
  • We introduce spade, our system that automatically generates data quality assertions for LLM pipelines. In a public release of spade’s candidate assertion generation component [^1] https://spade-beta.streamlit.app, we observe and analyze 2000+ deployments (Section 2.4),
  • We present a method to select a minimal set of assertions while meeting coverage and accuracy requirements, used by spade to reduce the number of assertions. For low-data settings, we introduce assertion subsumption as a novel proxy for coverage (Section 3), and
  • We demonstrate spade’s effectiveness on nine real-world LLM pipelines (eight of which we open-source). In our low-data setting (approximately 75 inputs and outputs per pipeline), our subsumption-based solution outperforms simpler baselines that do not consider interactions between assertions by reducing the number of assertions by 14% and lowering the false failure rate by 21% (Section 4).
2. Identifying Candidate Assertions
Figure 2. Taxonomy of prompt deltas from 19 LLM pipelines.
Version $i$ $\Delta\mathcal{P}_{\_}i$ $\Delta\mathcal{P}_{\_}i$ Category Possible New Assertion Criteria
1 + Given the following information about the user, {personal_info}, and information about a movie, {movie_info}: write a personalized note for why the user should watch this movie. Inclusion Response should be personalized and relevant to the given user information
2 + Include elements from the movie’s genre, cast, and themes that align with the user’s interests. Inclusion Response includes specific references to the user’s interests related to the movie’s genre, cast, and themes
3 + Ensure the recommendation note is concise. Qualitative Assessment Response should be concise
4 - Ensure the recommendation note is concise. + Ensure the recommendation note is concise, not exceeding 100 words. Count Response should be within the 100 word limit
5 - Include elements from the movie’s genre, cast, and themes that align with the user’s interests. + Mention the movie’s genre and any shared cast members between the {movie_name} and other movies the user has watched. Inclusion Response should mention genre and verify cast members are accurate
6 + Mention any awards or critical acclaim received by movie_name. Inclusion Response should include references to awards or critical acclaim of the movie
7 + Do not mention anything related to the user’s race, ethnicity, or any other sensitive attributes. Exclusion Response should not include references to sensitive personal attributes
Table 1. Comparison of 7 prompt versions for an LLM pipeline to write personalized movie recommendations. In each $\Delta\mathcal{P}_{\_}i$, a sentence starts with “+” if it is a newly added sentence; a sentence starts with “-” if it is removed from $\mathcal{P}_{\_}{i-1}$. For each $\Delta\mathcal{P}_{\_}i$, we list its category (using the taxonomy in Figure 2) and possible new assertion criteria.
Our first goal is to generate a set of candidate assertions. We describe how prompt deltas can inform candidate assertions and explain how to derive candidate assertions from them.
2.1. Prompt Deltas
A single-step LLM pipeline consists of a prompt template $\mathcal{P}$, which is formatted with a serialized input tuple $t$ to derive a prompt that is fed to an LLM, returning a response. There can be many versions of $\mathcal{P}$, depending on how a developer iterates on their prompt template. Let $\mathcal{P}_{\_}0$ be the empty string, the 0th version, and let $\mathcal{P}_{\_}i$ be the ith version of a template. In the movie recommendation example from the introduction, suppose there are 7 versions, where $\mathcal{P}_{\_}7$ is the following: “Given the following information about the user, {personal_info}, and information about a movie, {movie_info}: write a personalized note for why the user should watch this movie. Ensure the recommendation note is concise, not exceeding 100 words. Mention the movie’s genre and any shared cast members between the {movie_name} and other movies the user has watched. Mention any awards or critical acclaim received by {movie_name}. Do not mention anything related to the user’s race, ethnicity, or any other sensitive attributes.”
We define a prompt delta $\Delta\mathcal{P}_{\_}{i+1}$ to be the diff (or difference) between $\mathcal{P}_{\_}i$ and $\mathcal{P}_{\_}{i+1}$. Concretely, a prompt delta $\Delta\mathcal{P}$ is a set of sentences, where each sentence is tagged as an addition (i.e., “+”) or deletion (i.e., “-”). Table 1 shows the $\Delta\mathcal{P}$s for a number of versions for our example. Each sentence in $\Delta\mathcal{P}_{\_}i$ is composed of additions (i.e., new sentences in $\mathcal{P}_{\_}i$ that didn’t exist in $\mathcal{P}_{\_}{i-1}$) and deletions (i.e., sentences in $\mathcal{P}_{\_}{i-1}$ that don’t exist in $\mathcal{P}_{\_}{i}$). A modification to a sentence is represented by a deletion and addition—for example, $\Delta\mathcal{P}_{\_}{6}$ in Table 1 contains some new instructions added to a sentence from $\mathcal{P}_{\_}{5}$. Each addition in $\Delta\mathcal{P}_{\_}{i}$ indicates possible assertion criteria, as shown in the right-most column of Table 1.
2.2. Prompt Delta Analysis
To understand what assertions developers may care about, we turn to real-world LLM pipelines. We analyzed 19 LLM pipelines collected from LangChain users, each of which consists of between three and 11 historical prompt template versions. These pipelines span various tasks across more than five domains (e.g., finance, marketing, coding, education, health), from generating workout summaries to a chatbot acting as a statistics tutor. Table 7 in Appendix B shows a summary of the pipelines, including a description of each pipeline and the number of prompt versions. For each pipeline, we categorized prompt deltas, i.e., $\Delta\mathcal{P}_{\_}i$, into different types—for example, instructing the LLM to include a new phrase in each response (i.e., inclusion), or instructing the LLM to respond with a certain tone (i.e., qualitative criteria). Two authors iterated on the categories 4 times through a process of open and axial coding, ultimately producing the taxonomy in Figure 2. The taxonomy-annotated dataset of prompt versions can be found online [^2] https://github.com/shreyashankar/spade-experiments/blob/main/taxonomylabels.csv.
We divide deltas into two main high-level categories: Structural and Content-Based. Around 35% categories identified across all deltas in our dataset were Structural, and 65% were Content-Based. Structural deltas indicate a minor restructuring of the prompt, without changing any criteria of a good response (e.g., adding a newline for readability), or specification of the intended output (e.g., JSON or Markdown). Plausible assertion criteria based on structural deltas would check if the LLM output adheres to the user-specified structure. On the other hand, content-based deltas indicate a change in the meaning or definition of the task. Content-based deltas include descriptions of the workflow steps that the LLM should perform (e.g., “first, do X, then, come up with Y”), instructions of specific phrases to include or exclude in responses, or qualitative indicators of good responses (e.g., “maintain a professional tone”). The Data Integration subcategory (under Content-Based deltas) concerns adding new sources of context to the prompt—for example, adding a new variable like “{movie_info}” to the prompt, indicating a new type of information to be analyzed along with other content in the prompt. For some illustrative examples of prompt deltas for each category, we categorize the prompt deltas in Table 1, and in Table 2, we show sample prompt deltas for each category in our taxonomy. This taxonomy may be of independent interest to researchers studying the process of prompt engineering, as well as practitioners seeking to identify ways to improve their prompts for production LLM pipelines.
Overall, our exercise in building this taxonomy reveals two key findings. First, developers across diverse LLM pipelines iterate on prompts in similar ways as encoded as nodes in the taxonomy, many of which correspond to aspects that can be explicitly checked via assertions. Second, we find that an automated approach to synthesize data quality assertions may be promising, since our taxonomy reveals several such instances where the prompt delta could directly correspond to an assertion that captures the same requirement. For example, many deltas correspond to instructions to include or exclude specific phrases, indicating that developers may benefit from assertions that explicitly verify the inclusion or exclusion of such phrases in LLM outputs.
However, if we are to use an LLM to automatically generate assertion criteria based on our taxonomy of deltas, we also need to test whether LLMs can accurately identify the categories corresponding to a delta. Therefore, we confirmed GPT-4’s correct categorization of prompt deltas (as of October 2023): we assigned ground truth categories to all prompt deltas from the 19 pipelines, and GPT-4 achieved an F1 score of 0.8135. The prompt used for category extraction from prompt deltas is detailed in Appendix C.
Category Explanation Example Prompt Delta
Response Format Instruction Structure guidelines. “+ Start response with ‘You might like…’.”
Example Demonstration Illustrative example. “+ For example, here is a response for sci-fi fans…”
Prompt Clarification Refines prompt/removes ambiguity. “- Discuss/+ Explain movie fit…”
Workflow Description Describe “thinking” process. “+ First, analyze viewing history…”
Data Integration Adds placeholders. “+ Include user’s {genre} reviews.”
Quantity Instruction Adds numerical content. “+ Keep note under 100 words.”
Inclusion Instruction Directs specific content. “+ Mention movie awards/acclaim.”
Exclusion Instruction Advises on omissions. “+ Avoid movie plot spoilers.”
Qualitative Criteria Sets stylistic attributes. “+ Maintain friendly, positive tone.”
Table 2. Categories of prompt deltas.
2.3. From Taxonomy to Assertions
$\spadesuit$ spade Construct $\Delta\mathcal{P}_{\_}5$ “- Include elements from the movie’s genre, cast, and themes that align with the user’s interests. + Mention the movie’s genre and any shared cast members…” Prompt LLM with $\Delta\mathcal{P}_{\_}5$ to identify the delta type(s) and assertion criteria based on our taxonomy Synthesize Function(s)
Figure 3. Generating candidate assertions from a $\Delta\mathcal{P}$.
As we saw previously, prompt delta as categorized into nodes in our taxonomy often correspond to meaningful assertion criteria. Next, we need a method to automatically synthesize the data quality assertions from the prompt deltas. A natural idea is to prompt an LLM to generate assertion functions corresponding to relevant categories in our taxonomy given the prompt deltas. We tried this approach with GPT-4 in January 2024 and observed several omitted assertions that clearly corresponded to categories in our taxonomy. Therefore, we adopt a two-step prompting process in our approach, as it has been demonstrated that breaking tasks into steps can enhance LLM accuracy (Wei et al., 2022; Wu et al., 2022b; Grunde-McLaughlin et al., 2023). In the first step, we prompt GPT-4 for natural language descriptions of assertion criteria, and in the second step, we prompt GPT-4 to generate Python functions that implement such criteria. However, it’s worth noting that with daily advancements in LLMs, a one-step process might already be feasible. Additionally, while our taxonomy guides LLM-generated assertions now, future LLMs may implicitly learn these categories through reinforcement learning from human feedback (Christiano et al., 2017). Nevertheless, knowing the taxonomy-based categories associated with candidate assertions may help in filtering them.
More specifically, for each prompt delta $\Delta\mathcal{P}_{\_}i$, we first prompt an LLM to suggest as many criteria as possible for assertions—each aligning with a taxonomy category from Figure 3. A criterion is loosely defined as some natural language expression that operates on a given output or example and evaluates to True or False (e.g., “check for conciseness”). Our method analyzes every $\Delta\mathcal{P}_{\_}i$ instead of just the last prompt version for several reasons: developers often remove instructions from prompts to reduce costs while expecting the same behavior (Parnin et al., 2023), prompts contain inherent ambiguities and imply multiple ways of evaluating some criteria, and complex prompts may lead to missed assertions if only one version is analyzed. Consequently, analyzing each $\Delta\mathcal{P}_{\_}i$ increases the likelihood of generating relevant assertions.
For each delta, our method collects the criteria identified and prompts an LLM again to create Python assertion functions. The synthesized functions can use external Python libraries or pose binary queries to an LLM for complex criteria. For function synthesis, the LLM is instructed that if the criterion is vaguely specified or open to interpretation, such as “check for conciseness,” it can generate multiple functions that each evaluate the criterion. In this conciseness example, the LLM could return multiple functions—a function that splits the response into sentences and ensures that there are no more than, say, 3 sentences, a function that splits the response into words and ensures that there are no more than, say, 25 words, or a function that sends the response to an LLM and asks whether the response is concise. The overall outcome of this is a multiset of candidate functions $F=\{f_{\_}1,\dots,f_{\_}m\}$. The two prompts for generating assertion criteria and functions can be found in Appendix C.
2.4. Initial Deployment
To assess the potential of our auto-generated assertions, in November 2023, we released an early prototype of spade’s candidate assertion generation framework via a Streamlit application [^3] https://spade-beta.streamlit.app/. At the time of public release, the Streamlit application had a different set of prompts to generate assertions; these prompts generated relatively few assertions compared to the number of assertions generated in this paper’s version of spade. However, the taxonomy has remained the same. In the Streamlit app, a developer can either paste their prompt template that they want to generate assertions for, or they can point to their LangChain Hub prompt template (which contains prompt version histories via commits). The app then visualizes the identified taxonomy categories in the user’s prompt and displays the candidate assertions, as shown in the screenshot in Figure 4.
Tool Usage Insights and Feedback. From the reception to our Streamlit application, we found significant interest in auto-generated data quality assertions for LLM pipelines: there have been over 2000 runs of the app for custom prompt templates (i.e., not the sample default prompt template in app). These runs span many fields, including medicine, education, cooking, and finance—providing insights from a diversity of use cases for LLM pipelines. Figure 5 shows a rough breakdown of the use cases people wanted to generate assertions for; however, it’s important to note that some runs may not cleanly fit into a single category (e.g., a chatbot for telehealth-related questions for a medical provider could be in “customer support” and “health”). Interestingly, 8% of tasks related to conversational assistants, and we observed instances of at least four different companies generating assertions for their chatbots, given that the company name was in the prompt. Users for 48 of these runs clicked the “download assertions” button, which downloads the candidate assertions as a Python file. We note that users can also directly copy the assertions code displayed, instead of downloading the assertions as a Python file, and we did not measure the copy events. No users clicked the “thumbs down” button—which is not to say that the generated assertions were perfect. In fact, anecdotally, we have seen the candidate assertion quality improve over the last several months as GPT-4 continually improves. On average, 3.3 assertions were generated per run, with minimum 1 and maximum 10 assertions generated for a run. Most of the runs corresponded to a single prompt version. After our Streamlit app was released, we found an unprompted review of spade-generated assertions from a LangChain user who built a “chat-with-your-pdf” tool [^4] https://twitter.com/thcalafatidis/status/1728144652119769394: in their words, “When I saw it I didnt beleive it could work that well, but it really did and made the evaluation process fun and ez.”
Observations about Assertion Criteria. Across the 2000+ runs, with regards to our taxonomy, assertion criteria were most commonly derived from inclusion and exclusion instructions. For example, for a shopping assistant, where the objective was to find the most relevant product related to a customer’s query, responses were required to include “all the features [the customer] asked for.” In another customer support agent example, responses were required to include “exact quotes in the [context] relevant to the [customer’s] question…word for word.” One common exclusion instruction across chat-related tasks was to avoid any discussion unrelated to the end-user’s question; however, we noticed that GPT-4 struggled to generate assertion criteria around such a generic exclusion instruction. We found Python functions generated to implement criteria such as “avoid unrelated ideas”, did not use an LLM to check the criteria. Instead, they checked for the presence of specific phrases like “unrelated” and “I’m sorry, I did not find anything related” in the response. Such errors indicate the limitations of current state-of-the-art LLMs and may suggest the need for specialized, fine-tuned LLMs for generating assertions in future work; here, we work with the limitations of present-day state-of-the-art LLMs. We discuss this further in Section 4.4.
A new pattern we noticed after deployment, which wasn’t seen in our initial analysis, is that developers often wish to hide certain parts of the LLM workflow in the outputs. This could be viewed as a special instance of the exclusion category in our taxonomy. For instance, in a prompt related to enterprise automation, the first instruction was to write a query targeting a specific database (e.g., “Write a SQL query to fetch the most relevant table from the MySQL database”). However, another instruction in the prompt specified that the name of the database should not appear in the final summary returned to the end-user (e.g., “Do not mention that you queried the MySQL table X”). We were pleasantly surprised to find that our process for generating criteria successfully identified and implemented such exclusion instructions accurately.
Figure 4. Screenshot of an early version of the spade Streamlit application.
$\spadesuit$ spade 00 055 551010 10101515 15152020 2020 Health IT and Programming Project Management Customer Support Education Sports and Entertainment Cooking Finance Marketing Others 20.820.8 20.820.816.716.7 16.716.712.512.5 $12.5$ 12.512.58.38.3 $8.3$ $8.3$ 8.38.34.24.2 $4.2$ $4.2$ 4.24.2Percentage
Figure 5. Prompts submitted to the spade Streamlit application span a variety of fields.
Observations about Assertions. In LLM pipelines with numerous prompt versions, we observed two main patterns. First, prompt engineering often leads to many similar assertions, and redundancy can be a headache at deployment when a developer has to keep track of so many assertions. For instance, a pipeline to summarize lecture transcripts [^5] https://smith.langchain.com/hub/kirby/simple-lecture-summary had 14 prompt versions, with many edits localized to the paragraph providing instructions on titles, speakers, and dates, creating multiple overlapping assertions. 10 assertions assessed the lecture’s central thesis; two are as follows:
$\spadesuit$ spade Second, many assertions may be incorrect, causing runtime errors or false failures. Assessing the accuracy of these assertions is challenging, particularly for those that are complex or invoke LLMs themselves, where even experienced developers might not be able to gauge their effectiveness without viewing the results of the functions on many examples. Even if the assertions are not incorrect, they still may have undesirable failure rates or coverage, which is often hard for a developer to reason about in conjunction with other assertions. Since there can be 50+ assertions generated for just a handful of prompt versions (as demonstrated in Section 4), manually filtering them by eyeballing failure rates for each subset of assertions is impractical. Therefore, we adopt an automated approach for filtering, as discussed in the following section.
3. Filtering Candidate Assertions
As we saw in the previous section, we often have redundant and incorrect assertions, particularly in pipelines with numerous prompt versions. Here, we focus on filtering this candidate set to a smaller number, which not only improves efficiency and reduces cost when deploying the assertions to run in production, but it also reduces cognitive overhead for the developer.
3.1. Definitions
Consider $e_{\_}i$ as an example end-to-end execution (i.e., run) of an LLM pipeline on some input. For the purposes of this section, we assume the prompt template $\mathcal{P}$ is fixed to be the final prompt version that the developer eventually decided on. We denote $y_{\_}i\in\{0,1\}$ to represent whether the developer considers $e_{\_}i$ to be a success (1) or failure (0).
Let $E$ be the set of all such example runs (this set is not provided upfront, as we will deal with shortly). We define an assertion function $f:E\rightarrow\{0,1\}$, where 11 11 indicates that the example was processed successfully by the LLM and 00 0 otherwise. Let $F^{\prime}=\{f_{\_}1,f_{\_}2,\ldots,f_{\_}k\}$ be a set of $k$ data quality assertions. An example $e_{\_}i$ is deemed successful by $F^{\prime}$ if and only if it satisfies all assertions in $F^{\prime}$. Specifically:
$\hat{y}_{\_}i=\begin{cases}1~{}~{}\text{if }f(e_{\_}i)=1\quad\forall f\in F^{\prime},&\text{ ~{}~{}~{}~{} // Deemed successful by all of }F^{\prime}\\ 0~{}~{}\text{otherwise.}&\text{ ~{}~{}~{}~{} // Deemed a failure by}\geq 1\text{ of }F^{\prime}\end{cases}$
Given all $m$ candidate assertions $F=\{f_{\_}1,f_{\_}2,\ldots,f_{\_}m\}$, the objective is to select $F^{\prime}\subseteq F$ such that $\hat{y}_{\_}i=y_{\_}i$ for most examples in $E$, with $F^{\prime}$ being as small as possible. This goal involves maximizing failure coverage and minimizing the false failure rate and selected function count, expressed as follows:
Definition 3.1.
Coverage for a set $F^{\prime}$ is the proportion of actual failures that are correctly identified by $F^{\prime}$, defined as:
$\text{Coverage}\left(F^{\prime}\right)=\frac{\sum_{\_}{i}\mathbb{I}\left[\hat{y}_{\_}i=0\land y_{\_}i=0\right]}{\sum_{\_}{i}\mathbb{I}\left[y_{\_}i=0\right]}\text{ ~{}~{}~{}~{} // Failures caught by }F^{\prime}$
Definition 3.2.
False Failure Rate (FFR) for a set $F^{\prime}$ is the fraction of examples that $F^{\prime}$ incorrectly evaluates as failures ($\hat{y}_{\_}i=0$) when they are actually successful ($y_{\_}i=1$), defined as:
$\text{FFR}\left(F^{\prime}\right)=\frac{\sum_{\_}{i}\mathbb{I}\left[\hat{y}_{\_}i=0\land y_{\_}i=1\right]}{\sum_{\_}{i}\mathbb{I}\left[y_{\_}i=1\right]}\text{ ~{}~{}~{}~{} // Non-failures flagged by }F^{\prime}$
In both definitions above, $\hat{y}_{\_}i$ represents set $F^{\prime}$’s prediction for the $i$-th example, $y_{\_}i$ is the actual outcome, while $\mathbb{I}$ is the indicator function. In practice, coverage and FFR are impossible to compute since the universe of examples $E$ is unknown. So, for now, we assume access to a subset $E^{\prime}\subset E$ of labeled LLM responses, where $E^{\prime}$ is a manually provided set of example runs and may not contain all the types of failures the LLM pipeline could observe—an issue we will deal with in Section 3.3. Thus, we replace Definition 3.1 and Definition 3.2 with $\text{Coverage}_{\_}{E^{\prime}}\left(F^{\prime}\right)$ and $\text{FFR}_{\_}{E^{\prime}}\left(F^{\prime}\right)$, omitting the subscript $E^{\prime}$ for brevity. Table 8 in Appendix A summarizes the notation used throughout this section.
3.2. Coverage Problem Formulation
Our goal is therefore to select a minimal set of assertions $F^{\prime}\subseteq F$ based on a sample $E^{\prime}=\{e_{\_}1,\ldots,e_{\_}n\}\subset E$. Formally:
$\displaystyle\text{minimize}\quad|F^{\prime}|$
[…]
The above problem may be infeasible with certain values of $\alpha$, for multiple reasons. To see this, consider the case where no assertion catches a specific failure, while $\alpha$ is set to 11 11. We ignore this (and similar cases) for now and defer their discussion to Section 4. To expand out the above, we introduce a matrix $M$ (size $n\times m$) to track each assertion’s result on each example $e_{\_}i$, where $M_{\_}{ij}=1$ if $f_{\_}j(e_{\_}i)=1$ (i.e., $F_{\_}j$ deems $e_{\_}i$ a success) and $M_{\_}{ij}=0$ otherwise. We also define binary variables $x_{\_}j$ and $w_{\_}{ij}$ to represent whether an assertion is chosen and if it marks an example as a failure: $w_{\_}{ij}=\left(1-M_{\_}{ij}\right)\cdot x_{\_}j$, which is based on whether $f_{\_}j$ denotes $e_{\_}i$ as a failure and $f_{\_}j$ is included in $F^{\prime}$. We additionally introduce the binary variable $u_{\_}i$ to represent whether a failed example is covered by any selected assertion:
$\displaystyle u_{\_}i\leq\sum_{\_}{j=1}^{m}w_{\_}{ij},\quad\forall i\in\left[1,n\right]:y_{\_}i=0.\text{~{}~{}~{}~{}~{}~{}~{}~{}// Coverage of failure }e_{\_}i$
Then, the coverage constraint can be written as:
_$\displaystyle\frac{\sum_{\_}{i:y_{\_}i=0}u_{\_}i}{\sum_{\_}i\mathbb{I}\left[y_{\_}i=0\right]}\geq\alpha.$
Now, we formulate the FFR constraint. Observe that FFR can be decomposed into the following:
(1) _$\displaystyle\frac{\sum_{\_}{i=1}^{n}y_{\_}i\cdot\max_{\_}j\left(w_{\_}{ij}\right)}{\sum_{\_}{i=1}^{n}\left[y_{\_}i=1\right]}\leq\tau.$
The product $w_{\_}{ij}=\left(1-M_{\_}{ij}\right)\cdot x_{\_}j$ in the numerator is 1 if $f_{\_}j$ is included in $F^{\prime}$ and incorrectly marks a good example as a failure. The $\max$ function ensures that as long as there is some selected function $f_{\_}j$ that marks $e_{\_}i$ as a failure, the product is 1. Multiplying $y_{\_}i$ by the $\max$ result ensures that the numerator is incremented only when $y_{\_}i=1$, or when the example $e_{\_}i$ is actually not a failure. This numerator is thus equivalent to the numerator in Definition 3.2, i.e., $y_{\_}i=1\land\hat{y}_{\_}i=0$. The entire fraction is less than the threshold $\tau$, which represents the maximum allowed false failure rate across the examples. To formulate FFR according to (1), we introduce a new binary variable $z_{\_}i$, which defines whether $F^{\prime}$ denotes $e_{\_}i$ as a failure while $e_{\_}i$ is actually a successful example (i.e., false failure):
$\displaystyle z_{\_}i\geq y_{\_}i\cdot w_{\_}{ij},\quad\forall i\in\left[1,n\right];\,\forall j\in\left[1,m\right].\text{~{}~{}~{}~{}~{}~{}~{}~{}// }e_{\_}i\text{ is a false failure}$
Then, the FFR constraint is:
_$\displaystyle\frac{\sum_{\_}{i=1}^{n}z_{\_}i}{\sum_{\_}{i=1}^{n}\mathbb{I}\left[y_{\_}i=1\right]}\leq\tau.$
We can then state the problem of minimizing the number of assertions while meeting $E^{\prime}$ coverage and FFR constraints as an Integer Linear Program (ILP):
$\displaystyle\text{minimize}\quad\sum_{\_}{j=1}^{m}x_{\_}j$
[…]
$\displaystyle\quad u_{\_}i\leq\sum_{\_}{j=1}^{m}w_{\_}{ij},\quad\forall i\in\left[1,n\right]\ \text{where}\ y_{\_}i=0;\quad\frac{\sum_{\_}{i:y_{\_}i=0}u_{\_}i}{\sum_{\_}i\mathbb{I}\left[y_{\_}i=0\right]}\geq\alpha;$
$\displaystyle\quad z_{\_}i\geq y_{\_}i\cdot w_{\_}{ij},\quad\forall i\in\left[1,n\right],\ \forall j\in\left[1,m\right];\quad\frac{\sum_{\_}{i=1}^{n}z_{\_}i}{\sum_{\_}{i=1}^{n}\mathbb{I}\left[y_{\_}i=1\right]}\leq\tau;$
$\displaystyle\quad x_{\_}j,u_{\_}i,z_{\_}i,w_{\_}{ij}\in\{0,1\},\quad\forall i\in\left[1,n\right],\ \forall j\in\left[1,m\right].$
We refer to a solution for this ILP as $\textsc{spade}_{\_}{\text{cov}}$. Trivially, the problem is NP-hard for $\tau=0$ and $\alpha=1$, via a simple reduction from set cover, and is in NP, since it can be stated in ILP form. In our case, given tens of candidate assertions and fewer than 100 examples $e_{\_}i$, the ILPs tend to be of reasonable size (i.e., thousands of variables). Most ILP solvers can efficiently and quickly such programs.
3.3. Subsumption Problem Formulation
So far, we’ve assumed that the developer is willing to provide a comprehensive set of labeled example runs $E^{\prime}$. In settings where the developer is unwilling to do so, and where $E^{\prime}$ does not include all failure types in $E$, $\textsc{spade}_{\_}{\text{cov}}$ may overlook useful assertions in $F$ that only catch failures in $E\setminus E^{\prime}$—as shown empirically in Section 4. We initially considered using active learning (Castro and Nowak, 2008) to sample more LLM responses for each assertion and weak supervision to label the responses (Ratner et al., 2020). However, this approach can be costly with state-of-the-art LLMs, and it demands significant manual effort to balance failing and successful examples for each assertion, ensuring meaningful FFRs, and avoiding the exclusion of assertions due to underrepresented failure types. For this setting, we additionally introduce the notion of subsumption. Assuming that all candidate assertion functions cover as many failure modes as possible, our goal is to pick $F^{\prime}\subseteq F$ such that assertions in $F\setminus F^{\prime}$ are subsumed by $F^{\prime}$. Formally, a set of functions $S$ subsumes some function $f$ if the conjunction of functions in $S$ logically implies the conjunction of functions in $S$ and $f$. That is,
Definition 3.3.
A set of functions $S\implies f$ if and only if $\forall e\in E,\exists s\in S\text{ such that }s(e)\implies f(e)$. In other words, if $S\implies f$, then $f$ catches no new failures that $S$ does not already catch.
A simple example of subsumption is as follows: suppose our set of functions $S$ contains only one function $f$, which parses an LLM output into a JSON list to check that the list has at least 2 elements. $f$ may be generated as a result of a count-related instruction from our taxonomy of prompt deltas. Now, let $g$ be some other function that only checks if the output can be turned into a JSON list. $g$ might have been generated due to a “Response Format Instruction”-type delta. If $g$ fails for some LLM output, $f$—and therefore $S$—must also fail that output; thus $S\implies g$.
3.3.1. ILP with Subsumption Constraints
We reformulate the problem with subsumption constraints. Let $G$ be the set of functions in $F\setminus F^{\prime}$ not subsumed by $F^{\prime}$:
$\displaystyle\text{minimize}\quad|F^{\prime}|+|G|$
[…]
To represent $G$, we introduce binary variables and a matrix $K$ to denote subsumption relationships. $K_{\_}{ij}=1$ if and only if ${f_{\_}i}\implies f_{\_}j$. We discuss how we construct $K$ in more detail in Section 3.3.2. Recall that $x_{\_}j$ represents whether $f_{\_}j$ is selected in $F^{\prime}$. For each function $f_{\_}j$, a binary variable $r_{\_}j$ indicates if it is subsumed by $F^{\prime}$.
$\displaystyle x_{\_}i\cdot K_{\_}{ij}\leq r_{\_}j\leq\sum_{\_}{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{m}\left(x_{\_}{i}\cdot K_{\_}{ij}\right),\quad\forall j\in\left[1,m\right],i\neq j.$
Now, $s_{\_}j$ will denote if $f_{\_}j$ is neither in $F^{\prime}$ nor subsumed by $F^{\prime}$. We break this into three parts. First,
$\displaystyle s_{\_}j\leq 1-x_{\_}j,\quad\forall j\in\left[1,m\right],$
indicates that $f_{\_}j\not\in F^{\prime}$ if $s_{\_}j$ is allowed to be 1 (i.e., $f_{\_}j\not\in F^{\prime}$ when $x_{\_}j=0$). The second constraint,
$\displaystyle s_{\_}j\leq 1-r_{\_}j,\quad\forall j\in\left[1,m\right],$
captures the condition where $f_{\_}j$ is not subsumed by $F^{\prime}$. Here, if $r_{\_}j=0$, suggesting no subsumption, then $s_{\_}j$ may be 1. Lastly, to combine these conditions, we employ the constraint,
$\displaystyle s_{\_}j\geq 1-x_{\_}j-r_{\_}j,\quad\forall j\in\left[1,m\right],$
which ensures that $s_{\_}j$ can only be 1 if both prior conditions are satisfied: $f_{\_}j$ is neither in $F^{\prime}$ nor subsumed by $F^{\prime}$.
Finally, our objective is to minimize the sum of the number of functions in $F^{\prime}$ and non-subsumed functions $G$. The ILP formulation then becomes (with changes highlighted in blue):
$\displaystyle\text{minimize}\quad\sum_{\_}{j=1}^{m}x_{\_}j+{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}\sum_{\_}{j=1}^{m}s_{\_}j}$
[…]
$\displaystyle\quad u_{\_}i\leq\sum_{\_}{j=1}^{m}w_{\_}{ij},\quad\forall i\in\left[1,n\right]\ \text{where}\ y_{\_}i=0;\quad\frac{\sum_{\_}{i:y_{\_}i=0}u_{\_}i}{\sum_{\_}i\mathbb{I}\left[y_{\_}i=0\right]}\geq\alpha;$
$\displaystyle\quad z_{\_}i\geq y_{\_}i\cdot w_{\_}{ij},\quad\forall i\in\left[1,n\right],\forall j\in\left[1,m\right];\quad\frac{\sum_{\_}{i=1}^{n}z_{\_}i}{\sum_{\_}{i=1}^{n}\mathbb{I}\left[y_{\_}i=1\right]}\leq\tau;$
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}x_{\_}i\cdot K_{\_}{ij}\leq r_{\_}j\leq\sum_{\_}{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{m}\left(x_{\_}{i}\cdot K_{\_}{ij}\right),\quad\forall j\in\left[1,m\right],i\neq j;}$
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}s_{\_}j\leq 1-x_{\_}j,\quad s_{\_}j\leq 1-r_{\_}j,\quad\forall j\in\left[1,m\right];}$
$\displaystyle\quad{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}s_{\_}j\geq 1-x_{\_}j-r_{\_}j,\quad\forall j\in\left[1,m\right];}$
$\displaystyle\quad x_{\_}j,u_{\_}i,z_{\_}i,w_{\_}{ij},{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}r_{\_}j},{\color[rgb]{0,0,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.6875}s_{\_}j}\in\{0,1\},\quad\forall i\in\left[1,n\right],\ \forall j\in\left[1,m\right].$
We call a solution to this ILP $\textsc{spade}_{\_}{\text{sub}}$. We maintain the coverage constraint because the subsumption approach alone does not inherently account for the distribution or the significance of different types of failures. For instance, if a particular type of failure makes up a critical portion of $E^{\prime}$, a subsumption-based approach might overlook it. To see this in the simplest case, consider $\alpha=1$: simply optimizing for the sum $|F^{\prime}|+|G|$ does not guarantee all failures in $E^{\prime}$ are covered. In practice, $\textsc{spade}_{\_}{\text{sub}}$ is less sensitive to $\alpha$ than $\textsc{spade}_{\_}{\text{cov}}$, as we will discuss further in Section 4.
3.3.2. Assessing Subsumption
Here, we detail how to construct $K$, our matrix representing subsumption relationships between pairs of functions $f_{\_}i,f_{\_}j$. For pure Python functions, one could use static analysis to determine subsumption. However, it becomes complex when dealing with assertions that include LLM calls or a mix of pure Python and LLM-invoking assertions. For these, spade employs GPT-4 to identify potential subsumptions $\{a\}\implies b$ for pairs of functions $a,b$. For any pipeline, there are only two calls to the LLM to determine all subsumptions: first, all assertion functions are combined into a single prompt for GPT-4, instructing it to list as many subsumption relationships as it can identify, then prompting it again to transform its response into a parse-able list of pairs $a\implies b$. Details of this LLM subsumption prompt are in Appendix C.
To maximize precision of $\implies$ relationships identified, we employ some heuristics. First, $E^{\prime}$ can filter subsumptions: for $f_{\_}i$ and $f_{\_}j$, observe that:
[…]
In other words, any set containing $f_{\_}i$ definitely does not subsume $f_{\_}j$ if $f_{\_}j$ flags a failure that $\{f_{\_}i\}$ does not. Next, we use the FFR threshold to skip evaluating subsumption. Observe that, for any set of assertions $S$ and $f\not\in S$,
$\displaystyle\max\left(\text{FFR}\left(S\right),\text{FFR}\left(\{f\}\right)\right)$ $\displaystyle\leq\text{FFR}\left(S\cup\{f\}\right),$
(2) $\displaystyle\text{FFR}\left(S\cup\{f\}\right)$ $\displaystyle\leq\text{FFR}\left(S\right)+\text{FFR}\left(\{f\}\right).$
As such, we need not evaluate $\left\{f_{\_}i\right\}\implies f_{\_}j$ if either $\text{FFR}\left(\{f_{\_}i\}\right)\geq\tau$ or $\text{FFR}\left(\{f_{\_}j\}\right)\geq\tau$. Lastly, we use transitivity of implication to further prune checks: if ${x}\implies y$ and ${y}\implies z$, then ${x}\implies z$ is also true.
3.3.3. Subsumption without Examples
For completeness, we also consider the case where we have no developer-provided examples $E^{\prime}$. In this case, we are only reliant on our subsumption relationships to pick our set $F^{\prime}$. Our problem can then be restated as follows:
$\displaystyle\text{minimize}\quad\sum_{\_}{j=1}^{m}x_{\_}j+\sum_{\_}{j=1}^{m}s_{\_}j$
[…]
$\displaystyle\quad s_{\_}j\leq 1-x_{\_}j,\quad s_{\_}j\leq 1-r_{\_}j,\quad\forall j\in\left[1,m\right];$
$\displaystyle\quad s_{\_}j\geq 1-x_{\_}j-r_{\_}j,\quad\forall j\in\left[1,m\right];$
$\displaystyle\quad x_{\_}j,r_{\_}j,s_{\_}j\in\{0,1\},\quad\forall j\in\left[1,m\right].$
However, it turns out this problem is no longer NP-Hard. Consider the example in Figure 6, where an edge indicates subsumption. For example $a\rightarrow b$ means that $b$ catches a subset of the failures of $a$, and is therefore subsumed by $a$. Here, if we were to minimize $\Sigma_{\_}j(x_{\_}j+s_{\_}j)$, we would simply pick $a$ and $e$ to be part of $F^{\prime}$, since the rest of the functions would be subsumed (therefore their $s_{\_}j=0$—recall that $s_{\_}j=1$ for assertions that are neither subsumed nor selected), and our overall metric would evaluate to 22 22. One intuitive way to view this problem is to start by keeping all of the nodes as unselected, i.e., $x_{\_}j=0,s_{\_}j=1$, and then by adding them one at a time in a predefined order to $F^{\prime}$, we set $x_{\_}j=1,s_{\_}j=0$, and at the same time, we also impact $s_{\_}j$ (moving them from 11 11 to 00 0) for all other functions that then become subsumed as a result.
$\spadesuit$ spade abdcef
Figure 6. Example depicting subsumption relationships.
More generally, subsumption relationships can be represented in the form of a Directed Acyclic Graph (DAG), as in our example above. (We omit the trivial case where two or more functions are equivalent, that is the only case where there can be cycles; in all other cases we have a DAG.) Within this DAG, we simply select all nodes that have no incoming edges and add them to $F^{\prime}$. We can see that the rest of the nodes are subsumed, since they have at least one incoming edge. We could certainly exclude a node that doesn’t have any incoming edge from $F^{\prime}$, but adding it to $F^{\prime}$ does not worsen the objective because there is no way that that node will be subsumed by another. So, overall, we need to do a topological sort of the subsumption graph and pick the nodes “at the top”. Thus, the problem is in PTIME when there are no examples provided.
4. Evaluation
Pipeline # Good Ex. # Bad Ex. # Prompt Ver. Data Generation Task Full Prompt Link
codereviews 60 16 8 Writing reviews of GitHub repo pull requests homanp/github-code-reviews
emails 43 55 3 Creating SaaS user onboarding emails gitmaxd/onboard-email
fashion 48 34 16 Suggesting outfit ideas for specific events N/A
financea 48 52 5 Summarizing financial earnings call transcripts casazza/maptemplate
lecturesummariesb 27 22 14 Summarizing lectures or talks, focusing on main points and critical insights kirby/simple-lecture-summary
negotiation 27 19 8 Writing tailored negotiation strategies based on provided contracts and target prices antoniogonc/strategy-report
sportroutine 19 31 3 Transforming workout video transcripts into structured exercise routines aaalexlit/sport-routine-to-program
statsbot 39 31 3 Writing interactive discussions for any topic in statistics anthonynolan/statistics-teacher
threads 50 56 4 Crafting concise, engaging Twitter threads for specific audiences and topics flflo/summarization
Table 3. Description of data-generating LLM pipelines in our experiments. The fashion examples (and ground-truth indicators of whether the example response is good or bad) are provided by a startup that uses LangChain. All other examples are synthetically generated except examples for the finance and lecturesummaries pipelines, which are taken from Kaggle.
Pipeline # CA Method FFR Coverage on $E^{\prime}$ Frac Func. Selected Frac Excl. Funcs. not Subsumed
codereviews 44 baseline 0.117 1 0.456 (20) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0 0.625 0.045 (2) 0.409 (18)
$\textsc{spade}_{\_}{\text{sub}}$ 0.117 0.875 0.341 (15) 0 (0)
emails 24 baseline 0 1 0.5 (12) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0 1 0.0417 (1) 0.458 (11)
$\textsc{spade}_{\_}{\text{sub}}$ 0 1 0.458 (11) 0 (0)
fashion 106 baseline 0.878 0.971 0.632 (67) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0.245 0.6 0.028 (3) 0.321 (34)
$\textsc{spade}_{\_}{\text{sub}}$ 0.224 0.62 0.377 (40) 0 (0)
finance 47 baseline 0.667 1 0.787 (37) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0.229 0.673 0.085 (4) 0.553 (26)
$\textsc{spade}_{\_}{\text{sub}}$ 0.208 0.981 0.553 (26) 0 (0)
lecturesum. 70 baseline 0.528 1 0.457 (32) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0.194 0.643 0.014 (1) 0.414 (29)
$\textsc{spade}_{\_}{\text{sub}}$ 0.194 1 0.343 (24) 0 (0)
negotiation 50 baseline 0.444 1 0.4 (20) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0.222 0.632 0.04 (2) 0.32 (16)
$\textsc{spade}_{\_}{\text{sub}}$ 0.185 1 0.34 (17) 0 (0)
sportroutine 26 baseline 0.211 1 0.538 (14) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0.211 0.774 0.077 (2) 0.462 (12)
$\textsc{spade}_{\_}{\text{sub}}$ 0 0.871 0.308 (8) 0 (0)
statsbot 15 baseline 0 1 0.467 (7) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0 0.935 0.133 (2) 0.333 (5)
$\textsc{spade}_{\_}{\text{sub}}$ 0 1 0.467 (7) 0 (0)
threads 34 baseline 0 1 0.765 (26) 0 (0)
$\textsc{spade}_{\_}{\text{cov}}$ 0 0.875 0.029 (1) 0.735 (25)
$\textsc{spade}_{\_}{\text{sub}}$ 0 1 0.589 (20) 0 (0)
Table 4. Results of different versions of spade with $\alpha=0.6$ and $\tau=0.25$. “# CA” is short for the number of candidate assertions. The ✓and ✗ marks denote whether $\alpha$ and $\tau$ constraints are met. Each entry is a fraction of the total number of candidate assertions for that pipeline (with the absolute number in parentheses). $\textsc{spade}_{\_}{\text{cov}}$ selects the fewest assertions overall. $\textsc{spade}_{\_}{\text{sub}}$ selects the fewest assertions while optimizing for subsumption.
We first discuss the LLM pipelines and datasets (i.e., $E^{\prime}$); then, we discuss methods and metrics and present our results. The experiment code, datasets, and LLM responses are hosted on GitHub [^6] https://github.com/shreyashankar/spade-experiments.
4.1. Pipeline and Dataset Descriptions
We evaluated spade on nine LLM pipelines—eight from LangChain Hub [^7] https://smith.langchain.com/hub, an open-source collection of LLM pipelines, and 1 proprietary pipeline. Each pipeline consists of a prompt template for the LLM and collection of approximately 75 examples (i.e., inputs and ouptuts in $E^{\prime}$) with labels for whether the outputs were good or bad. All nine pipelines come from LangChain users. Six pipelines were used in developing our prompt delta taxonomy (Section 2.2), each representing a different domain (e.g., programming, finance, marketing). Two pipelines came from spade’s Streamlit deployment (Section 2.4). We include one final proprietary pipeline (the fashion pipeline) in our experiments because it had the largest number of prompt template versions and spade’s assertions have already been deployed in production for tens of thousands of daily runs.
Table 3 provides details on each LLM pipeline and corresponding set of good and bad examples. For the fashion pipeline, good and bad examples were provided by a developer at the corresponding startup. While we used real user-provided prompt templates and histories (between 3 and 16 prompt versions) for the other 8 pipelines, we constructed and annotated our own input-output examples so we could release them publicly. For two pipelines, we sourced examples from Kaggle. For the other six pipelines, we synthetically generated the other datasets using Chat GPT Pro (based on GPT-4) and manually reviewed them. For instance, for the codereviews pipeline that uses an LLM to review pull requests, we asked Chat GPT to generate example pull requests covering a variety of programming languages, application types, and diff sizes. On average, we collected 75 examples per pipeline. We then executed the LLM pipelines on these inputs and manually labeled the responses to assess whether they met the prompt instructions.
4.2. Method Comparison and Metrics
As before, let $E^{\prime}$ be a dataset of example prompt-response pairs, as well as the corresponding labels of whether the response was good (i.e., 1) or bad (i.e., 0). Let $\tau$ be the FFR threshold and $F$ be the set of candidate assertions produced by the first step of spade (Section 2). If a candidate function $f$ results in a runtime error for some example $e$, we denote $f\left(e\right)=0$ (i.e., failure). All of our code was written in Python, using the PuLP Python package to find solutions for the ILPs. We used the default PuLP configuration, which uses the CBC solver (Forrest and Lougee-Heimer, 2005).
The simplest baseline involves generating candidate assertions and choosing all of them, but it proved ineffective, yielding 100% coverage and 100% FFR due to at least one assertion failing all tests. Failures were due to errors or overly specific conditions that, in theory, could pass some outputs (for example, requiring a precise phrase in a certain case) but, in reality, never matched any actual LLM outputs. As such, we evaluated two versions of spade against a baseline that simply filters candidate assertions that individually exceed the FFR threshold:
  • baseline selects all functions $f$ in $F$ where $\text{FFR}\left(\{f\}\right)\leq\tau$
  • $\textsc{spade}_{\_}\text{cov}$ is a solution to the ILP defined in Section 3.2
  • $\textsc{spade}_{\_}\text{sub}$ is a solution to the ILP defined in Section 3.3.1
Let $F^{\prime}$ represent the set of selected assertions by any version of spade. We measure four metrics:
  1. (1)
    Fraction of Assertions Selected (i.e., $|F^{\prime}|/|F|$)
  2. (2)
[…]
  3. (3)
    False Failure Rate (Definition 3.2)
  4. (4)
    Coverage on $E^{\prime}$ (Definition 3.1)
Additionally, an important aspect of $\textsc{spade}_{\_}{\text{sub}}$’s success is the effectiveness of subsumption assessment between all pairs of assertions. Since we do not have ground truth for subsumption, we focus on precision, calculated as the proportion of correctly identified subsumed pairs out of all subsumed pairs identified by the LLM. We do not assess recall—whether GPT-4 identified every possible subsumption—due to the impracticality of labeling possibly tens of thousands of assertion pairs per pipeline. Moreover, precision is more critical than recall or accuracy, as identifying even some subsumptions allows $\textsc{spade}_{\_}{\text{sub}}$ to achieve a solution with fewer selected assertions than baseline.
4.3. Results and Discussion
Pipeline Precision
codereviews 0.90
emails 0.79
fashion 0.74
finance 0.79
lecturesummaries 0.89
negotiation 0.68
sportroutine 0.89
statsbot 0.86
threads 0.80
Table 5. Precision of assessing subsumption with GPT-4 (verified by two authors).
Pipeline $\textsc{spade}_{\_}{\text{cov}}$ $\textsc{spade}_{\_}{\text{sub}}$
codereviews 0.267 0.362
emails 0.196 0.225
fashion 0.628 1.197
finance 0.352 0.441
lecturesummaries 0.265 0.538
negotiation 0.220 0.332
sportroutine 0.120 0.158
statsbot 0.104 0.122
threads 0.272 0.332
Table 6. Average ILP runtimes (in seconds) over 10 trials for $\textsc{spade}_{\_}{\text{cov}}$ and $\textsc{spade}_{\_}{\text{sub}}$. The baseline method has no ILP component.
First, we briefly discuss whether individual components of spade are practical (e.g., determining subsumption, solving the ILPs). Using GPT-4 to assess subsumption results in an average precision of 0.82 across all pipelines, as seen in Table 5, confirming its effectiveness. The ILP runtimes in Table 6 are all under one second, except for 1.197 seconds for $\textsc{spade}_{\_}{\text{sub}}$ for the fashion pipeline. This relatively small runtime demonstrates the feasibility of our approach. In the remainder of this section, we will focus on the assertions chosen by different methods.
For simplicity, we set the coverage and FFR thresholds to be the same across all pipelines ($\alpha=0.6$, $\tau=0.25$). We report results for the three methods in Table 4. Consider the codereviews pipeline, for example, which uses an LLM to review a pull request for any code repository. Here, baseline selects 20 assertions, $\textsc{spade}_{\_}{\text{cov}}$ selects two assertions, and $\textsc{spade}_{\_}{\text{sub}}$ selects 15 assertions. By selecting more functions, $\textsc{spade}_{\_}{\text{sub}}$ ensures that all non-subsumed functions are included. All three approaches respect the $E^{\prime}$ coverage constraint, but baseline violates the FFR constraint in 4 out of 9 pipelines, while the spade approaches do not violate the FFR constraint.
For our workloads, $\textsc{spade}_{\_}{\text{sub}}$ opts for approximately 14% fewer assertions compared to baseline and shows a significantly lower FFR, reducing it by about 21% relative to baseline. $\textsc{spade}_{\_}{\text{cov}}$ excludes, on average, about 44% of functions that are not subsumed by $F^{\prime}$. Choosing a spade implementation primarily depends on how much labeled data is available. We subsequently discuss the trade-offs between different spade implementations.
Subsumption vs. $E^{\prime}$ Coverage. $\textsc{spade}_{\_}{\text{cov}}$ and $\textsc{spade}_{\_}{\text{sub}}$ are complementary, the former being more useful if $E^{\prime}$ is more comprehensive. For our datasets, $E^{\prime}$ is not comprehensive: Table 4 reveals that, on average, 44% of functions excluded by $\textsc{spade}_{\_}{\text{cov}}$ are not subsumed by the selected functions, despite being accurate within the FFR threshold. This is unsurprising given that each task has only 34 bad (i.e., failure) examples on average. While larger or more mature organizations may have extensive datasets and could get a meaningful result from $\textsc{spade}_{\_}{\text{cov}}$, $\textsc{spade}_{\_}{\text{sub}}$’s ability to select assertions that cover unrepresented potential failures can be beneficial in data-scarce settings. For example, here is a sample of 3 assertions for the codereviews pipeline ignored by $\textsc{spade}_{\_}{\text{cov}}$ but included in $\textsc{spade}_{\_}{\text{sub}}$ (with comments excluded for brevity):
$\spadesuit$ spade Subsumption as a Means for Reducing Redundancy. Several pipelines exhibit a large discrepancy between functions selected in baseline and $\textsc{spade}_{\_}{\text{sub}}$, which occurs when there are many redundant candidates. For example, in the codereviews pipeline’s 8 prompt versions, the developer iterated several times on the instruction to give a clear and concise review, resulting in five assertions that check the same thing (two of which are shown below): $\spadesuit$ spade Since all the five assertions meet the FFR constraint, individually, baseline would select them all, which is undesirable because they all do the same thing, but $\textsc{spade}_{\_}{\text{sub}}$ would select the one most compatible with the FFR constraint, as long as subsumption is assessed correctly. On the flip side, while assessing subsumption, the LLM may not recall all subsumptions, so $\textsc{spade}_{\_}{\text{sub}}$ may have duplicate assertions. For example, the codereviews pipeline contains assertions titled assert_includes_code_improvementv1 and assert_includes_code_improvementv2. $\boldsymbol{\alpha}$ and $\boldsymbol{\tau}$ Threshold Sensitivity. The feasibility of solutions from the ILP solver in spade is dependent on the chosen $\alpha$ and $\tau$ thresholds. If a feasible solution is not found, developers may need to adjust these values in a binary search fashion. In our case, all 9 LLM pipelines yielded feasible solutions with $\alpha=0.6$ and $\tau=0.25$. However, the small size of $E^{\prime}$ makes $\textsc{spade}_{\_}{\text{cov}}$ particularly sensitive to $\alpha$. In the pipelines, we observed that between one and five assertions covered 60% of $E^{\prime}$’s failures. For example, $\textsc{spade}_{\_}{\text{cov}}$ selected only one assertion for the emails pipeline: $\spadesuit$ spade If $E^{\prime}$ is exhaustive of failure modes and representative of the distribution of failures (e.g., for the emails pipeline, most failures are actually due to the response lacking an encouragement to contact the company), $\textsc{spade}_{\_}{\text{cov}}$ might be a satisfactory solution. However, our $E^{\prime}$ datasets clearly were not exhaustive, considering that $\textsc{spade}_{\_}{\text{sub}}$ always chose additional assertions. $\textsc{spade}_{\_}{\text{sub}}$ is less sensitive to $\alpha$, as it explicitly selects assertions based on their potential to cover new failures (i.e., subsumption) without exceeding the FFR, even if the constraint on coverage is no longer tight. FFR Tradeoffs. Considering that the difference between the fraction of functions selected for baseline and $\textsc{spade}_{\_}{\text{sub}}$ is less than 10% for three LLM pipelines, one may wonder if the complexity of $\textsc{spade}_{\_}{\text{sub}}$ is worth it. $\textsc{spade}_{\_}{\text{sub}}$ is generally preferable because baseline fails to consistently meet the FFR threshold $\tau$. We observed that as prompt versions increase, so do the number of assertions, impacting baseline adversely. The worst-case FFR of a set is the sum of individual FFRs, as shown in Equation 2. Hence, with a large number of independent assertions, the total FFR is likely to surpass the threshold. This issue is evident in the fashion and lecturesummaries pipelines, where despite each of the 67 and 32 assertions meeting FFR constraints individually, the total FFR for baseline reaches 88% and 53%, respectively. In practice, if spade were to be deployed in an interactive system, where spade could observe each LLM call in real-time (e.g., as a wrapper around the OpenAI API), the multitude of prompt versions further necessitates filtering assertions based on overall FFR. This underscores the need for the more complex $\textsc{spade}_{\_}{\text{cov}}$ or $\textsc{spade}_{\_}{\text{sub}}$ approaches.
4.4. Limitations and Future Work
Improving Quality of LLMs. While LLMs (both closed and open-source) are improving rapidly, and we don’t explicitly study prompt engineering strategies for spade, a complementary research idea is to explore such strategies or fine-tune small open-source models to generate assertions. Moreover, we proposed subsumption as a coverage proxy but didn’t explore prompt engineering or even non-LLM strategies (e.g., assertion provenance) for assessing subsumption. Despite LLM advancements, spade’s filtering stage remains crucial for reducing redundancy and ensuring accuracy, especially since assertions may involve LLMs.
Collecting Labeled Examples. Acquiring labeled data ($E^{\prime}$) is hard. Most of our datasets had very few prompt versions (only the ones committed to LangChain Hub), but in reality, developers may iterate on their prompt tens or hundreds of times. Future work could involve passive example collection via LLM API wrappers or gathering developer feedback on assertions. Prioritizing different types of assertions and formalizing these priorities within spade is another area for exploration. Additionally, assessing the accuracy of FFR estimates with limited $E^{\prime}$ and exploring methods to enhance FFR accuracy in the absence of large labeled datasets (e.g., via prediction-powered inference (Angelopoulos et al., 2023)), presents an interesting area for future work.
Supporting More Complex LLM Pipelines. Our study focuses on single-prompt LLM pipelines, but more complex pipelines, such as those involving multiple prompts, external resources, and human interaction, present opportunities for auto-generating assertions for each pipeline component. For instance, in retrieval-augmented generation pipelines(Lewis et al., 2020), assertions could be applied to the retrieved context before LLM responses are even generated.
5. Related Work
We survey work from prompt engineering, evaluating ML and LLMs, LLMs for software testing, and testing ML pipelines.
Prompt Engineering. For both nontechnical (Zamfirescu-Pereira et al., 2023) and technical users (Si et al., 2022; Parameswaran et al., 2023), prompt engineering is hard for several reasons: small changes in prompt phrasing (Liu et al., 2023; Arora et al., 2022) or the order of instructions or contexts (Lu et al., 2021) can significantly affect outputs. Moreover, as LLMs change under the hood of the API (i.e., prompt drift), outputs can change without developer awareness (Chen et al., 2023). Tools and papers are emerging to aid in prompt management and experimentation, and are even using LLMs to write prompts (Arawjo et al., 2023; Wu et al., 2022b, a; Cheng et al., 2023; Zhou et al., 2022). Moreover, deployed prompts introduce new challenges, like “balancing more context with fewer tokens” and “wrangling prompt output” to meet user-defined criteria (Parnin et al., 2023). Our work doesn’t focus explicitly on helping developers create better prompts, but it could indirectly support developers by helping identify mistakes.
ML and LLM Evaluation. Evaluating and monitoring deployed ML models is known to be challenging (Paleyes et al., 2022; Shankar et al., 2022b). Evaluating LLMs in a deployed setting is even more challenging because LLMs are typically used for generative tasks, where the outputs are free-form (Dou et al., 2021). Some LLM pipeline types, like question-answering with retrieval-augmented generation pipelines (Lewis et al., 2020), can use standardized, automated metrics (Es et al., 2023; Saad-Falcon et al., 2023), but others face challenges due to unknown metrics and lack of labeled datasets (Xiao et al., 2023; Chang et al., 2023; Parnin et al., 2023). Typically, organizations rely on human evaluators for LLM outputs (Gehrmann et al., 2023; Wang et al., 2023c; Parnin et al., 2023), but recent studies suggest LLMs can self-evaluate effectively with detailed “scorecards” (Chan et al., 2023; Zhang et al., 2023; Kim et al., 2023). However, writing these scorecards can be challenging (Parnin et al., 2023), motivating auto-generated evaluators. Recent work (Rebedea et al., 2023; Singhvi et al., 2023; Kim et al., 2023) and industry tools (Guardrails, 2023; Langchain, 2023; Llama Index, 2023) proposes the use of assertions to catch mistakes in LLM pipelines, while requiring the user to select these assertions.
LLMs for Software Testing. LLMs are increasingly being used in software testing, mainly for generating unit tests and test cases (Wang et al., 2023a; Schäfer et al., 2023; Siddiq et al., 2023; Steenhoek et al., 2023; Lemieux et al., 2023). Research explores how LLMs’ prompting strategies, hallucinations, and nondeterminism affect code or test accuracy (Ouyang et al., 2023; Feldt et al., 2023; Chen et al., 2021; Fan et al., 2023). Our work is complementary and leverages LLMs to generate code-based assertions for LLM pipelines.
Testing and Validation in ML Pipelines. ML pipelines are hard to manage in production. Much of the literature on ML testing is geared towards validating structured data, through analyzing data quality (Hynes et al., 2017; Breck et al., 2019b; Schelter et al., 2018; Shankar et al., 2023a) or provenance (Schelter et al., 2023; Namaki et al., 2020). Platforms for ML testing typically offer automated experiment tracking and prevention against overfitting (Renggli et al., 2019; Agrawal et al., 2019), as well as tools for data distribution debugging (Grafberger et al., 2022). Model-specific assertions typically require human specification (Kang et al., 2020), or at least large amounts of data to train learned assertions (Kang et al., 2022). LLM chains or pipelines are a new class of ML pipelines, and LLMs themselves can generate assertions with little data. A recent study highlights the difficulty of testing LLM pipelines for “copilot”-like products: developers want to ensure accuracy while avoiding excessive resource use, such as running hundreds of assertions (Parnin et al., 2023)—motivating assertion filtering.
6. Conclusion and Future Work
We introduce a new problem of auto-generating assertions to catch failures in LLM pipelines, as well as spade, our framework for doing so. spade comprises two components: first, it synthesizes candidate assertions, and then it filters them down into a more manageable subset. To synthesize candidate assertions, we analyzed prompt version histories and learned that prompt deltas are often a rich source of requirements and therefore candidate assertions. We developed a taxonomy of prompt deltas for assertion synthesis, demonstrating its value via integration and deployment with LangChain, with over 2000 runs across domains. For the latter problem of candidate assertion filtering, we expressed the selection of an optimal set of assertions, covering most failures while introducing as few false failures as possible as an Integer Linear Program (ILP). We proposed assertion subsumption to cover failures in data-scarce scenarios and incorporated this into our ILP. We also studied the setting where there are no examples and demonstrated that it reduces to a topological sort on the subsumption graph. Our auto-generating assertion system, spade, was evaluated on nine real-world data-generating LLM pipelines. We have made our code and datasets public for further research and analysis.
There are a number of open questions in our effort to make production LLM pipelines more robust. For instance, while meeting developer provided criteria ($\alpha,\tau$) is helpful, sometimes developers would like to examine the generated and selected assertions in a way that helps them make the tradeoffs themselves, motivating a human-in-the-loop interface to assist developers in defining data quality for LLM pipelines. Such an interface could also be a vehicle for getting developers to label examples on the fly. Determining which labeled examples would help best select from the set of assertions is an open question that is reminiscent of active learning. Finally, we could also envision automatically updating assertion sets in deployed pipelines, as new failure modes inevitably arise in production.
Appendix A Notation
Table 8 summarizes the important notation used in the paper, including high-level descriptions of the ILP variables.
Symbol Description
$\mathcal{P}$ A prompt template (i.e., a string with placeholders, intended to be submitted to an LLM)
$\mathcal{P}_{\_}i$ The $i$th version of a prompt template
$\Delta\mathcal{P}_{\_}i$ The diff or changes between consecutive versions of a prompt template
$E$ The set of all hypothetical example runs for an LLM pipeline
$E^{\prime}$ A small set of example runs for an LLM pipeline (with labels denoting whether responses were good) that we can observe, where $E^{\prime}\subset E$
$n$ The number of examples in $E^{\prime}$
$e_{\_}i$ The $i$th example in $E^{\prime}$
$f$ An arbitrary assertion function
$F$ The set of all candidate assertions from Section 2
$m$ The number of assertions in $F$
$F^{\prime}$ A subset of $F$ selected as a minimal set of assertions
$G$ Set of functions in $F\setminus F^{\prime}$ not subsumed by $F^{\prime}$ (as defined by Definition 3.3)
$\hat{y}_{\_}i$ Binary indicator of whether example $e_{\_}i$ satisfies all assertions in $F^{\prime}$
$y_{\_}i$ Binary indicator of whether a developer considers the LLM pipeline response for example $e_{\_}i$ good (i.e., success) or bad (i.e., failure)
$\alpha$ Threshold for coverage of failing examples; $F^{\prime}$ should cover at least $\alpha$ of failing examples
$\tau$ Threshold for False Failure Rate (FFR); $F^{\prime}$ should fail no more than $\tau$ good examples
$M$ $n\times m$ matrix representing the results of assertions on examples in $E^{\prime}$
$K$ $m\times m$ matrix representing subsumption relationships between functions
$x_{\_}j$ ILP variable indicating inclusion of $f_{\_}j$ in $F^{\prime}$
$w_{\_}{ij}$ ILP variable representing if $F^{\prime}$ denotes $e_{\_}i$ as a failure
$u_{\_}i$ ILP variable indicating if a failed example is covered
$z_{\_}i$ ILP variable defining if $F^{\prime}$ incorrectly marks a successful example as a failure (false failure)
$r_{\_}j$ ILP variable indicating if a function is subsumed by any function in $F^{\prime}$
$s_{\_}j$ ILP variable representing functions in $F\setminus F^{\prime}$ not subsumed by $F^{\prime}$
Table 8. Notation used in the paper
Appendix B LangChain Prompt Hub and Taxonomy of Prompt Deltas
The LangChain Prompt Hub is an open-source repository of prompts for chains, detailing the version history of the prompts. In analyzing prompt deltas, we filter the Prompt Hub for user-uploaded chains with at least 3 prompt versions. Table 7 summarizes the chains analyzed.
Appendix C spade Prompts
spade leverages LLMs in three places. The first uses GPT-4 to categorize the delta (constructed by Python’s difflib). The second uses GPT-4 to generate Python assertion functions, based on the prompt and categories identified by the first ste.. The third usage of LLMs is in asking GPT-4 to evaluate whether two functions subsume each other.
Given prompt_diff, a list of sentences that have been modified from the previous prompt template to the current prompt template (i.e., $\Delta\mathcal{P}$), the prompt for categorizing the delta is as follows:
$\spadesuit$ spade Let concepts be the parsed categories identified by the previous prompt. The prompt for generating the Python assertion functions is as follows: $\spadesuit$ spade
C.1. Evaluating Subsumption
To evaluate subsumption, we use one prompt to query the subsumed pairs, given all assertions, and then a second prompt to format the subsumed pairs as a JSON so it can be easily parsed by spade. The first prompt is as follows:
$\spadesuit$ spade In the prompt above, the assertion_blob represents a string of all assertion functions. Then, the second prompt is as follows: $\spadesuit$ spade